### v1

# A New Algorithm for Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons

Abstract
We propose a new algorithm for Zero-Sum Linear-Quadratic Stochastic Differential Games by dissecting their inherent structures. Specifically, we construct dual-layer iterative matrix-increasing sequences, which reformulate the original problem into a set of mutually interconnected subproblems. By sequentially computing the stabilizing solutions to the associated algebraic Riccati equations within each subproblem, we derive the stabilizing solutions for the original problem and rigorously establish the convergence of the proposed algorithm. Numerical simulations further validate the effectiveness of the method. This work extends classical setting and provides the first complete unified numerical framework for solving a broader class of stochastic Game-Theoretic algebraic Riccati equations.
Keywords: Iteration Algorithm; Zero-Sum Linear-Quadratic Stochastic Differential Games; Game-Theoretic algebraic Riccati equations; Stabilizing Solutions
MSCcodes: 65F45, 91A15, 93D15
1 Introduction
Zero-Sum Stochastic Differential games provide a rigorous mathematical framework for modeling adversarial decision-making in stochastic dynamic systems. Building on the foundational work of [17] and [24] in differential games, [15] extended the deterministic framework developed by [20] to stochastic settings. Within this framework, the Zero-Sum Linear-Quadratic Stochastic Differential Game (ZSLQSDG) serves as a pivotal simplified model. It circumvents the significant challenges associated with solving the highly nonlinear Hamilton-Jacobi-Bellman-Isaacs equation, making it a key vehicle for both theoretical exploration and applied research. To date, significant advancements have been achieved in the theoretical study of ZSLQSDG; see, for example, [23, 27, 32, 31, 26, 34, 33, 29].
This paper focuses on the classic infinite-horizon ZSLQSDG with constant coefficients. For infinite-horizon problems–whether in deterministic or stochastic cases–only the stabilizing solution is of practical interest among all possible solutions. Specifically for this problem, [31] proves that the existence of closed-loop saddle points is characterized by the solvability of an algebraic Riccati equation (AREs) with an indefinite quadratic term under a certain stabilizing condition. In other words, solving a ZSLQSDG problem is equivalent to solving its associated AREs.
In the existing literature, AREs featuring indefinite quadratic terms are commonly referred to as Game-Theoretic algebraic Riccati equations (GTAREs). For further relevant studies, see [6, 25, 10, 3, 2, 12, 5, 33] and their references. Due to the indefiniteness of their quadratic terms, the traditional numerical algorithms for solving conventional AREs are no longer applicable. For deterministic GTAREs, [21] proposed a recursive method for the positive stabilizing solution. Unlike deterministic systems, stochastic systems with state-dependent noise complicate stabilizing solution computation. This complexity renders traditional deterministic methods inapplicable to the above cases. To address this, [14] developed an algorithm for GTAREs arising from a specific LQZSDG with state-dependent noise. This approach was later extended by [11] to stochastic $H_{\infty}$ problems involving state- and control-dependent noise, and further by [7] to stochastic $H_{\infty}$ problems involving state-, control-, and disturbance-dependent noise. [19] investigated a class GTAREs arising in stochastic $H_{\infty}$ problems with state-dependent noise. The study explored three effective approaches for computing stabilizing solutions of these equations, and further performed numerical comparisons and analyses of the iterative methods among them. Further computational studies can be found in [4, 9, 18, 8, 1, 30] and their references.
To date, no iterative algorithm has been reported in the existing literature for classical ZSLQSDG with state- and both-player-dependent noise. This paper develops a new algorithm for ZSLQSDG by dissecting their inherent structures. we construct dual-layer matrix iteration sequences via defect correction method proposed in [22], which reformulate the ZSLQSDG problem into a series of interconnected bilevel subproblems. These subproblems are solved sequentially by identifying the maximal solutions to their associated AREs, ultimately yielding stabilizing solutions for the original problem. Importantly, this study first extends the classical framework and develops a unified algorithm applicable to three classes of problems: deterministic GTAREs, stochastic $H_{\infty}$-type GTAREs, and the more general stochastic GTAREs associated with the classical ZSLQSDG.
Remaining Content Outline: Section 2 covers the mathematical framework of classical ZSLQSDG, a decoupled representation method, and the algorithm iteration process. Section 3 presents the main results, including the construction logic of the inner-outer matrix iteration sequence, the proofs of existence, uniqueness, and boundedness of the stabilizing solutions within each subproblem, as well as the analysis of algorithm convergence. Section 4 details the content related to numerical experiments.
2 Preliminary
2.1 Notation
Let us first introduce the following notation that will be used below:
  • $\mathbb{R}^{n}$: $n$-dimensional Euclidean space; $\mathbb{C}^{-}$: complex numbers with negative real part; $\mathbb{R}^{n\times m}$: the set of all $n\times m$ real matrices; $\mathbb{S}^{n}$: the set of all $n\times n$ symmetric matrices; $\overline{\mathbb{S}}_{+}^{n}$: the set of all $n\times n$ symmetric positive semi-definite matrices; $\mathbb{S}_{+}^{n}$: the set of all $n\times n$ symmetric positive definite matrices.
  • $I_{n}$: the identity matrix of size $n$; $\mathbb{O}_{n\times m}$: the null matrix of size $n\times m$. It can be simplified as 0 when no ambiguity is generated; $\mathbb{O}_{m\times n}^{\times(l)}$ denotes the object represented by $\mathbb{O}_{m\times n}$ arranged repeatedly for $l$ times.
  • $A^{\top}$: the transpose of the matrix $A$; $A^{\dagger}$: the Moore-Penrose pseudoinverse of the matrix $A$; $\mathrm{Tr}(\cdot)$: the trace of a square matrix; $\langle\cdot,\cdot\rangle$: the inner product on a Hilbert space. In particular, the usual inner product on $\mathbb{R}^{n\times m}$ is given by $\langle A,B\rangle\mapsto\mathrm{Tr}(A^{\top}B)$;
  • If $A\in\mathbb{S}_{+}^{n}$ (resp., $A\in\overline{\mathbb{S}}_{+}^{n}$) is a symmetric positive definite (resp., symmetric positive semi-definite) matrix, we write $A\succ 0$ (resp., $A\succeq 0$). For any $A,B\in\mathbb{S}^{n}$, we use the notation $A\succ B$ (resp., $A\succeq B$) to indicate that $A-B\succ 0$ (resp., $A-B\succeq 0$).
[…]
$L_{\mathbb{F}}^{2}(\mathbb{H})=\left\{\varphi:[0,\infty)\times\Omega\rightarrow\mathbb{H}\mid\varphi\in\mathbb{F}\text{ and }\mathbb{E}\int_{0}^{\infty}|\varphi(t)|^{2}\,dt<\infty\right\},\quad\mathcal{U}_{i}=L_{\mathbb{F}}^{2}(\mathbb{R}^{m_{i}}),\,i=1,2.$
$\mathcal{X}_{\mathrm{loc}}[0,\infty)=\left\{\varphi:[0,\infty)\times\Omega\rightarrow\mathbb{R}^{n}\mid\varphi\in\mathbb{F}\text{ is continuous, and }\mathbb{E}\left[\sup_{0\leq t\leq T}|\varphi(t)|^{2}\right]<\infty\text{ for every }T>0\right\},$
$\mathcal{X}[0,\infty)=\left\{\varphi\in\mathcal{X}_{\mathrm{loc}}[0,\infty)\mid\mathbb{E}\int_{0}^{\infty}|\varphi(t)|^{2}\,dt<\infty\right\}.$
2.2 Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a complete filtered probability space on which a r-dimensional standard Brownian motion $W=\{w_{1}(t),\dots,w_{r}(t);t\geq 0\}$ is defined with $\mathbb{F}=\{\mathcal{F}_{t}\}_{t\geq 0}$ being the usual augmentation of the natural filtration generated by $W$. We consider the following controlled linear stochastic differential equation (SDE) on $0,\infty)$:
$\begin{cases}dX(t)=[AX(t)+B_{1}u_{1}(t)+B_{2}u_{2}(t)]dt+\sum_{l=1}^{r}[C_{l}X(t)+D_{l,1}u_{1}(t)+D_{l,2}u_{2}(t)]dw_{l}(t)\\ X(0)=x\end{cases},$ (1)
in which $x\in\mathbb{R}^{n}$, $A,C_{l}\in\mathbb{R}^{n\times n}$ and $B_{i},D_{l,i}\in\mathbb{R}^{n\times m_{i}}$ ($i=1,2;1\leq l\leq r$). In the above, the state process $X$ is an $n$-dimensional vector, player 1 $u_{1}$ and player 2 $u_{2}$ are an $m_{1}$-dimensional vector and an $m_{2}$-dimensional vector, respectively.
In this paper, Player 1 and Player 2 share the same performance functional:
$J(x;u_{1},u_{2})\triangleq\mathbb{E}\int_{0}^{\infty}\Bigg[\Bigg\langle\begin{pmatrix}Q&S_{1}^{\top}&S_{2}^{\top}\\ S_{1}&R_{11}&R_{12}\\ S_{2}&R_{21}&R_{22}\end{pmatrix}\begin{pmatrix}X(t)\\ u_{1}(t)\\ u_{2}(t)\end{pmatrix},\begin{pmatrix}X(t)\\ u_{1}(t)\\ u_{2}(t)\end{pmatrix}\Bigg\rangle\Bigg]dt,$ (2)
the weighting coefficients in (2) satisfy:
$Q\in\mathbb{S}^{n},\quad R_{21}^{\top}=R_{12}\in\mathbb{R}^{m_{1}\times m_{2}},\quad S_{i}\in\mathbb{R}^{m_{i}\times n},R_{ii}\in\mathbb{S}^{m_{i}}(i=1,2).$
For $(x,u_{1},u_{2})\in\mathbb{R}^{n}\times\mathcal{U}_{1}\times\mathcal{U}_{2}$, the solution $X(\cdot;x,u_{1},u_{2})$ to the SDE in (1) may only exist in $\mathcal{X}_{loc}[0,\infty)$, which renders $J(x;u_{1},u_{2})$ ill-defined. We define the set of admissible controls as $\mathcal{U}_{ad}(x)\triangleq\{(u_{1},u_{2})\in\mathcal{U}_{1}\times\mathcal{U}_{2}\mid X(\cdot;x,u_{1},u_{2})\in\mathcal{X}[0,\infty)\}$. A pair $(u_{1},u_{2})\in\mathcal{U}_{ad}(x)$ is called an admissible control pair for the initial state $x$, and the corresponding $X(\cdot;x,u_{1},u_{2})$ is referred to as the admissible state process. In this case, $J(x,u_{1},u_{2})$ is clearly well-defined.
In this zero-sum game, Player 1 (the maximizer) selects control $u_{1}$ to maximize (2), while Player 2 (the minimizer) chooses $u_{2}$ to minimize the same function. The problem is to find an admissible control pair $(u_{1}^{*},u_{2}^{*})$ that both players can accept. We denote the above-mentioned problem as $({SDG})^{\,0}_{\infty}$ for short. For a description of the $({SDG})^{\,0}_{\infty}$ problem, refer to [28] . More detailed information can be found therein.
The $({SDG})^{\,0}_{\infty}$ problem corresponds to following stochastic GTAREs:
$\begin{cases}Q(P)-S(P)^{\top}R(P)^{\dagger}S(P)=0\\ \mathcal{R}(S(P))\subseteq\mathcal{R}(R(P))\\ R_{11}(P)\succeq 0,\quad R_{22}(P)\preceq 0\end{cases},$ (3)
where $\mathcal{R}(M)$ denotes the range of a matrix or an operator $M$,
$Q(P)=PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q,$
$S(P)=\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}=\begin{bmatrix}B_{1}^{\top}P+\sum_{l=1}^{r}D_{l,1}^{\top}PC_{l}+S_{1}\\ B_{2}^{\top}P+\sum_{l=1}^{r}D_{l,2}^{\top}PC_{l}+S_{2}\\ \end{bmatrix}$,
$R(P)=\begin{bmatrix}R_{11}(P)&R_{12}(P)\\ R_{21}(P)&R_{22}(P)\\ \end{bmatrix},\quad R_{ij}(P)=R_{ij}+\sum_{l=1}^{r}D_{l,i}^{\top}PD_{l,j}\,(i,j=1,2).$
Remark 2.1.
For the stochastic GTAREs (3), the general theoretical framework of $({SDG})^{\,0}_{\infty}$ only requires that $R(P)$ satisfies $\mathcal{R}(S(P))\subseteq\mathcal{R}(R(P))$ in [28]. However, in our subsequent discussion, we impose the slightly stronger condition that $R(P)$ be of full rank. This requirement is well-motivated as it allows us to focus on the core structural properties of the ZSLQSDG. We therefore restrict our analysis to the case where $R(P)$ is full rank.
Definition 2.2.
The system (1) donated as $(A,C_{1},\dotsb,C_{r})$ is called mean-square stable such that the solution of the following SDE:
$\begin{cases}dX(t)=AX(t)dt+\sum_{l=1}^{r}C_{l}X(t)dw_{l}(t)\\ X(0)=x\end{cases},$ (4)
satisfies $\lim_{t\to\infty}\mathbb{E}[X(t)^{\top}X(t)]=0$ for every initial state $x\in\mathbb{R}^{n}$.
A concept related to the above admissible control pair is that of mean-square stabilizers. We define the set of all such stabilizers associated with the system (1) as $\mathcal{K}=$
$\left\{(\varTheta_{1},\varTheta_{2})\in\mathbb{R}^{m_{1}\times n}\times\mathbb{R}^{m_{2}\times n}\ \middle|\ \begin{aligned} &\text{The system}\\ &(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2},C_{1}+D_{1,1}\varTheta_{1}+D_{1,2}\varTheta_{2},\dotsb,C_{r}+D_{r,1}\varTheta_{1}+D_{r,2}\varTheta_{2})\\ &\text{ is mean-square stable}.\end{aligned}\right\},$
Let $(\varTheta_{1},\varTheta_{2})\in\mathcal{K}$, the feedback control $U(\cdot)=\begin{bmatrix}\varTheta_{1}X(\cdot)\\ \varTheta_{2}X(\cdot)\end{bmatrix}$ is called mean-square stabilizing. We refer to [25, 28] for the above definition.
Definition 2.3 ([*]Sun2020_book).
A matrix $P\in\mathbb{S}^{n}$ is called a stabilizing solution of stochastic GTARE (3) if $P$ satisfies (3) and $(K_{1}(P),K_{2}(P))\in\mathcal{K}$, where $\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{\top}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}$ and $\mathcal{K}$ is the set of all such stabilizers associated with the system (1).
Remark 2.4.
Theorem 2.6.7 in [28] is a direct conclusion pertaining to the relationship between $({SDG})^{\,0}_{\infty}$ and stochastic GTAREs. Leveraging this theoretical result, we convert $({SDG})^{\,0}_{\infty}$ into an equivalent problem of finding the stabilizing solution $P$ to the stochastic GTARE (3).
To facilitate subsequent analysis of the problem, we introduce the following decoupling representation method. Setting
$\begin{bmatrix}v_{1}(t)\\ v_{2}(t)\end{bmatrix}=\begin{bmatrix}u_{1}(t)\\ u_{2}(t)\end{bmatrix}-\begin{bmatrix}K_{1}(0)\\ K_{2}(0)\end{bmatrix},\quad\begin{bmatrix}K_{1}(0)\\ K_{2}(0)\end{bmatrix}=-R(0)^{-1}\begin{bmatrix}S_{1}(0)\\ S_{2}(0)\end{bmatrix}$,
and $v_{2}(t)=Lx(t)$ in $({SDG})^{\,0}_{\infty}$, we obtain
$\begin{cases}dX(t)=[A_{L}X(t)+B_{1}v_{1}(t)]dt+\sum_{l=1}^{r}[C_{lL}X(t)+D_{l,1}v_{1}(t)]dw_{l}(t)\\ x(0)=x\end{cases}$ (5)
in which $x\in\mathbb{R}^{n}$, and
$J_{L}(x;v_{1})\triangleq\mathbb{E}\int_{0}^{\infty}\Bigg[\Bigg\langle\begin{pmatrix}Q_{L}&S^{\top}_{L}\\ S_{L}&R_{11}\end{pmatrix}\begin{pmatrix}X(t)\\ v_{1}(t)\end{pmatrix},\begin{pmatrix}X(t)\\ v_{1}(t)\end{pmatrix}\Bigg\rangle\Bigg]dt,$ (6)
where
$\begin{cases}A_{L}=A+B_{1}K_{1}(0)+B_{2}K_{2}(0)+B_{2}L,\,C_{lL}=C_{l}+D_{l,1}K_{1}(0)+D_{l,2}K_{2}(0)+D_{l,2}L,&1\leq l\leq r\\ Q_{L}=Q-S^{\top}(0)R(0)^{-1}S(0)+L^{\top}R_{22}L,\quad S_{L}=R_{12}L\end{cases}.$
The corresponding ARE is
$PA_{L}+A_{L}^{\top}P+\sum_{l=1}^{r}C_{lL}^{\top}PC_{lL}+Q_{L}$ (7)
$-\left(B_{1}^{\top}P+\sum_{l=1}^{r}D_{l,1}^{\top}PC_{lL}+S_{L}\right)^{\top}\left(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}PD_{l,1}\right)^{-1}\left(B_{1}^{\top}P+\sum_{l=1}^{r}D_{l,1}^{\top}PC_{lL}+S_{L}\right)=0.$
Throughout this work, $\mathcal{A}$ stands for the set of $L$, where $L\in\mathbb{R}^{m_{2}\times n}$ satisfy:
  • The system $(A_{L},C_{1L},\dotsb,C_{rL})$ is mean-square stable.
  • The corresponding ARE (7) has a stabilizing solution $\tilde{P}_{L}$, satisfying the sign conditions
    $R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1}\prec 0.$
2.3 Algorithm Design
In this paper, we aim to solve for the closed-loop saddle points numerically for the $({SDG})^{\,0}_{\infty}$ problem with given parameter values. Consider the sequences $\{P^{(k)}\}_{k\geq 0}$ and $\{Z^{(k)}\}_{k\geq 0}$ constructed according to the following procedure:
  1. 1.
    For $k=0$, set $P^{(0)}=0$, and let $Z^{(0)}$ be the unique stabilizing solution to the following algebraic Riccati equation with a definite quadratic term:
    $0=Z^{(0)}A_{(0)}+A_{(0)}^{\top}Z^{(0)}+\sum_{l=1}^{r}C_{l,(0)}^{\top}Z^{(0)}C_{l,(0)}+M_{(0)}$ (8)
    $-\left(B_{2}^{\top}Z^{(0)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}C_{l,(0)}\right)^{\top}\left(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}D_{l,2}\right)^{-1}\left(B_{2}^{\top}Z^{(0)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}C_{l,(0)}\right),$
    where $A_{(0)}=A-\begin{bmatrix}B_{1}&B_{2}\end{bmatrix}R(0)^{-1}S(0),\,C_{l,(0)}=C_{l}-\begin{bmatrix}D_{l,1}&D_{l,2}\end{bmatrix}R(0)^{-1}S(0),1\leq l\leq r$$M_{(0)}=Q-S^{\top}(0)R(0)^{-1}S(0)$.
  2. 2.
    For $k\geq 1$, update $P^{(k)}$ via:
    $P^{(k)}=P^{(k-1)}+Z^{(k-1)},$ (9)
    and solving $Z^{(k)}$, which is the unique stabilizing solution to the following algebraic Riccati equation with a definite quadratic term:
    $0=Z^{(k)}A_{(k)}+A_{(k)}^{\top}Z^{(k)}+\sum_{l=1}^{r}C_{l,(k)}^{\top}Z^{(k)}C_{l,(k)}+M_{(k)}$ (10)
    $-\left(B_{2}^{\top}Z^{(k)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}C_{l,(k)}\right)^{\top}\left(R_{22}(P^{(k)})+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}D_{l,2}\right)^{-1}\left(B_{2}^{\top}Z^{(k)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}C_{l,(k)}\right).$
    the matrices $A_{(k)},C_{l,(k)}$ and $M_{(k)}$ evolve as:
    $\begin{cases}A_{(k)}=A_{(k-1)}-\begin{bmatrix}B_{1}&B_{2}\end{bmatrix}R(P^{(k)})^{-1}\begin{bmatrix}N_{1,(k-1)}\\ N_{2,(k-1)}\end{bmatrix}\\ C_{l,(k)}=C_{l,(k-1)}-\begin{bmatrix}D_{l1}&D_{l2}\end{bmatrix}R(P^{(k)})^{-1}\begin{bmatrix}N_{1,(k-1)}\\ N_{2,(k-1)}\end{bmatrix},\quad 1\leq l\leq r\\ M_{(k)}=[N_{1,(k-1)}-R_{12}(P^{(k)})R_{22}(P^{(k)})^{-1}N_{2,(k-1)}]^{\top}R^{\sharp}_{22}(P^{(k)})^{-1}\\ \quad\quad\quad\times[N_{1,(k-1)}-R_{12}(P^{(k)})R_{22}(P^{(k)})^{-1}N_{2,(k-1)}]\end{cases}$ (11)
    where
    $R^{\sharp}_{22}(P)=R_{11}(P)-R_{12}(P)R_{22}(P)^{-1}R_{21}(P),\begin{bmatrix}N_{1,(k-1)}\\ N_{2,(k-1)}\end{bmatrix}=\begin{bmatrix}B_{1}^{\top}Z^{(k-1)}+\sum_{l=1}^{r}D_{l,1}^{\top}Z^{(k-1)}C_{l,(k-1)}\\ B_{2}^{\top}Z^{(k-1)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k-1)}C_{l,(k-1)}\end{bmatrix}$. (12)
Under some natural assumptions, we will show:
  • The sequences $\{Z^{(k)}\}$ and $\{P^{(k)}\}$ are well-defined. Each algebraic Riccati equation admits a unique stabilizing solution $\tilde{P}_{L}\succeq Z^{(k)}\succeq 0(k=0,1,2,3\cdots)$, where $\tilde{P}_{L}$ be the stabilizing solution of the ARE (7).
  • These sequences are convergent and we have
[…]
3 The Main Results
3.1 Analysis of Structural Characteristics for ZSLQSDG
From stochastic GTAREs (3), we can define a mapping $\mathcal{G}:\mathrm{Dom}\,\mathcal{G}\to\mathbb{S}^{n}$ that satisfies the following relation:
$\mathcal{G}(P)=PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q-S(P)^{\top}R(P)^{-1}S(P).$ (13)
The nonlinear function $\mathcal{G}$ is not defined over the entire space $\mathbb{S}^{n}$. Instead, it is well-defined on the subset:
$\mathrm{Dom}\,\mathcal{G}=\left\{P\in\mathbb{S}^{n}|R_{22}(P)\succ 0\,\,\text{and}\,\,R_{11}(P)\prec 0\right\}.$
The aforementioned mapping $\mathcal{G}$ serves as the fundamental foundation for constructing the algorithm. We begin by detailing its general properties, which collectively ensure the algorithm’s ability to operate iteratively.
Proposition 3.1.
Let $P$ and $Z$ be such that $P,P+Z\in\mathrm{Dom}\,\mathcal{G}$. Then the feedback gains satisfy the following relation:
$\begin{bmatrix}K_{1}(P+Z)\\ K_{2}(P+Z)\end{bmatrix}=\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}-R(P+Z)^{-1}\begin{bmatrix}B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\\ B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\end{bmatrix}$,
where
$\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}$.
Proof.
Define the increment,
$\Delta=\begin{bmatrix}K_{1}(P+Z)\\ K_{2}(P+Z)\end{bmatrix}-\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}$.
By leveraging the definitions of $K_{1}(\cdot),K_{2}(\cdot)$, and multiplying both sides of the equation by $R(P+Z)$, we obtain:
$R(P+Z)\Delta=R(P+Z)R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}-\begin{bmatrix}S_{1}(P+Z)\\ S_{2}(P+Z)\end{bmatrix}$.
Note that,
$R(P+Z)=R(P)+\Delta R,\Delta R=\begin{bmatrix}\sum_{l=1}^{r}D_{l,1}^{\top}ZD_{l,1}&\sum_{l=1}^{r}D_{l,1}^{\top}ZD_{l,2}\\ \sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,1}&\sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,2}\end{bmatrix}$.
Thus,
$R(P+Z)\Delta=\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}+\Delta R\cdot R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}-\begin{bmatrix}S_{1}(P+Z)\\ S_{2}(P+Z)\end{bmatrix}$.
Using the identity,
$R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}=-\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}$,
we obtain:
$R(P+Z)\Delta=-\begin{bmatrix}B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\\ B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\end{bmatrix}$.
Multiplying both sides by $R(P+Z)^{-1}$ yields:
$\begin{bmatrix}K_{1}(P+Z)\\ K_{2}(P+Z)\end{bmatrix}-\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P+Z)^{-1}\begin{bmatrix}B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\\ B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\end{bmatrix}$.
Proposition 3.2 ([*]Dragan2013book).
For any $P\in\mathrm{Dom}\,\mathcal{G}$ and for all $\varTheta_{1}\in\mathbb{R}^{m_{1}\times n},\varTheta_{2}\in\mathbb{R}^{m_{2}\times n}$, we have
$\mathcal{G}(P)={}$ $P(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2})+(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2})^{\top}P$
$+\sum_{l=1}^{r}(C_{l}+D_{l,1}\varTheta_{1}+D_{l,2}\varTheta_{2})^{\top}P(C_{l}+D_{l,1}\varTheta_{1}+D_{l,2}\varTheta_{2})+Q$
$+\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}^{\top}R\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}+\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}^{\top}\begin{bmatrix}S_{1}\\ S_{2}\end{bmatrix}+\begin{bmatrix}S_{1}\\ S_{2}\end{bmatrix}^{\top}\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}$
$-\begin{bmatrix}K_{1}(P)-\varTheta_{1}\\ K_{2}(P)-\varTheta_{2}\end{bmatrix}^{\top}R(P)\begin{bmatrix}K_{1}(P)-\varTheta_{1}\\ K_{2}(P)-\varTheta_{2}\end{bmatrix},\,\text{where}\,\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}$.
Proposition 3.3.
Let $P$, $Z$ satisfy $P,P+Z\in\mathrm{Dom}\,\mathcal{G}$. Denote the feedback gain matrix as:
$K(P)\triangleq\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}$,
and the auxiliary matrix as:
$N(P,Z)\triangleq\begin{bmatrix}N_{1}(P,Z)\\ N_{2}(P,Z)\end{bmatrix}=\begin{bmatrix}B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))\\ B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))\end{bmatrix}$. (14)
Then the following identity holds:
$\mathcal{G}(P+Z)=\mathcal{G}(P)+Z(A+B_{1}K_{1}(P)+B_{2}K_{2}(P))+(A+B_{1}K_{1}(P)+B_{2}K_{2}(P))^{\top}Z$
$+\sum_{l=1}^{r}(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))^{\top}Z(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))-N(P,Z)^{\top}R(P+Z)^{-1}N(P,Z).$
Proof.
Starting from the definition of $\mathcal{G}$,
$\mathcal{G}(P+Z)=(P+Z)A+A^{\top}(P+Z)+\sum_{l=1}^{r}C_{l}^{\top}(P+Z)C_{l}+Q-S(P+Z)^{\top}R(P+Z)^{-1}S(P+Z).$
From Proposition 3.2, expanding and rearranging terms, we obtain:
$\mathcal{G}(P+Z)=$ $PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q+K(P)^{\top}R(P)K(P)+S(P)^{\top}K(P)+K(P)^{\top}S(P)$
$+Z(A+B_{1}K_{1}(P)+B_{2}K_{2}(P))+(A+B_{1}K_{1}(P)+B_{2}K_{2}(P))^{\top}Z$
$+\sum_{l=1}^{r}(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))^{\top}Z(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))$
$-(K(P+Z)-K(P))^{\top}R(P+Z)(K(P+Z)-K(P)).$
From Proposition 3.1, we derive:
$(K(P+Z)-K(P))^{\top}R(P+Z)(K(P+Z)-K(P))=N(P,Z)^{\top}R(P+Z)^{-1}N(P,Z).$
Thus,
$\mathcal{G}(P+Z)=$ $\mathcal{G}(P)+Z(A+B_{1}K_{1}(P)+B_{2}K_{2}(P))+(A+B_{1}K_{1}(P)+B_{2}K_{2}(P))^{\top}Z$
$+\sum_{l=1}^{r}(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))^{\top}Z(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))$
$-N(P,Z)^{\top}R(P+Z)^{-1}N(P,Z).$
3.2 A Criteria for the Existence and Uniqueness of Stabilizing Solutions of AREs with Definite Quadratic Term
Based on the structure of $\mathcal{G}$ and Proposition 3.3, a series of interconnected sub-algebraic Riccati equations with definite quadratic terms can be constructed, as described in Section 2.3. A key requirement for the algorithm is the existence and uniqueness of stabilizing solutions to these equations. We present a criterion for this purpose in what follows.
Definition 3.4 ([*]Dragan2013book).
Given the following stochastic observation system:
$\begin{cases}dX(t)=AX(t)dt+\sum_{l=1}^{r}C_{l}X(t)dw_{l}(t)\\ dY(t)=E_{0}X(t)dt+\sum_{l=1}^{r}E_{l}X(t)dw_{l}(t)\end{cases},$ (15)
where $E_{l}\in R^{q\times n}$ for $0\leq l\leq r$, we denote this system by $\left[E_{0},E_{1},\dots,E_{r};A,C_{1},\dots,C_{r}\right]$. It is said to be stochastically detectable if there exists a constant matrix $\varTheta\in\mathbb{R}^{n\times q}$ such that the system
$\left(A+\varTheta E_{0},\ C_{1}+\varTheta E_{1},\ \dots,\ C_{r}+\varTheta E_{r}\right)$
is mean-square stable.
Lemma 3.5.
If the system $\left[E_{0},E_{1},\dots,E_{r};A,C_{1},\dots,C_{r}\right]$ is stochastically detectable, then the following statements are equivalent:
  • (a.)
    The system $(A,C_{1},\dots,C_{r})$ is mean-square stable. Let $\mathcal{L}^{*}$ denote the linear operator associated with this system, defined by
    $\mathcal{L}^{*}(P)=PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l},\quad\forall P\in\mathbb{S}^{n}.$
    Then all eigenvalues of $\mathcal{L}^{*}$ lie in the left half-plane, i.e., $\operatorname{Spec}\mathcal{L}^{*}\subset\mathbb{C}^{-}$.
  • (b.)
    There exists a matrix $P\in\overline{\mathbb{S}}_{+}^{n}$ satisfying the matrix equation:
    $PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+\sum_{l=0}^{r}E_{l}^{\top}E_{l}=0.$
Proof.
Proof of (a.) $\Rightarrow$ (b.): Since $\sum_{l=0}^{r}E_{l}^{\top}E_{l}\succeq 0$ and the system $(A,C_{1},\dots,C_{r})$ is mean-square stable, it follows from Theorem 3.2.2. and Theorem 2.7.5 in [12] that the equation
$PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+\sum_{l=0}^{r}E_{l}^{\top}E_{l}=0$
has a solution $P\in\overline{\mathbb{S}}_{+}^{n}$.
Proof of (b.) $\Rightarrow$ (a.): This direction follows directly from Remark 4.1.5 in [12]. For a detailed argument, see Theorem 4.1.7 and Remark 4.1.5 in the same reference. Furthermore, if the system $(A,C_{1},\dots,C_{r})$ is mean-square stable, then $\operatorname{Spec}\mathcal{L}^{*}\subset\mathbb{C}^{-}$; this is a consequence of Theorem 3.2.2. and Theorem 2.7.7 in [12]. ∎
Now we consider AREs of the form:
$PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q-\left(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}\right)^{\top}\left(R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\right)^{-1}\left(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}\right)=0,$ (16)
where $B,D_{l}\in\mathbb{R}^{n\times m}$ for $1\leq l\leq r$, and $R\in\mathbb{R}^{m\times m}$. The remaining parameters follow the previous definitions.
Define the operator $\bm{\Lambda}^{[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r};Q,R]}:\mathbb{S}^{n}\rightarrow\mathbb{S}^{n+m}$ associated with the ARE (16) as:
$\bm{\Lambda}^{[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r};Q,R]}(P)=\begin{bmatrix}PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q&\Big(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}\Big)^{\top}\\ B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}&R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\end{bmatrix}$. (17)
Define the set $\bm{\Gamma}^{[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r};Q,R]}$ related to ARE (16) as:
$\bm{\Gamma}^{[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r};Q,R]}=\left\{P\in\mathbb{S}^{n}\,\middle|\,\bm{\Lambda}(P)\succeq 0,\ R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\succ 0\right\}.$ (18)
Definition 3.6 ([*]Dragan2013book).
The system (16) is denoted by $\left[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r}\right]$ is said to be stochastically stabilizable if there exists a constant matrix $\varTheta\in\mathbb{R}^{n\times m}$ such that the system
$\left(A+B\varTheta,C_{1}+D_{1}\varTheta,\dots,C_{r}+D_{r}\varTheta\right)$
is mean-square stable.
Proposition 3.7.
Suppose the parameters of the ARE (16) satisfy the following:
  • $0\in\bm{\Gamma}^{[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r};Q,R,S]}$.
  • The system $A,C_{1},\dots,C_{r};\ B,D_{1},\dots,D_{r}]$ is mean-square stabilizable.
  • There exists a matrix set $\{E_{0},E_{1},\dots,E_{r}\}$ satisfying $\sum_{l=0}^{r}E_{l}^{\top}E_{l}=Q$ such that the system
    $E_{0},E_{1},\dots,E_{r};A,C_{1},\dots,C_{r}]$
    is stochastically detectable.
Then the ARE (16) admits a unique stabilizing solution $P\in\overline{\mathbb{S}}_{+}^{n}$ such that $R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\succ 0$.
Proof.
By Theorem 4.7 in [13], along with the condition $0\in\bm{\Gamma}$ and the mean-square stabilizability of $\left[A,C_{1},\dots,C_{r};B,D_{1},\dots,D_{r}\right],$ the ARE (16) has a maximal solution $P_{\text{max}}\in\overline{\mathbb{S}}_{+}^{n}$ such that $R+\sum_{l=1}^{r}D_{l}^{\top}P_{\text{max}}D_{l}\succ 0$. Using Proposition 3.2, the ARE for $P_{\text{max}}$ can be rewritten as:
$P_{\text{max}}\left(A+BT(P_{\text{max}})\right)+\left(A+BT(P_{\text{max}})\right)^{\top}P_{\text{max}}$
$+\sum_{l=1}^{r}\left(C_{l}+D_{l}T(P_{\text{max}})\right)^{\top}P_{\text{max}}\left(C_{l}+D_{l}T(P_{\text{max}})\right)+\sum_{l=0}^{r}\hat{E}_{l}^{\top}\hat{E}_{l}=0,$
where $T(P)=-\left(R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\right)^{-1}\left(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}\right)$, and
$\hat{E}_{l}=\begin{bmatrix}\mathbb{O}_{q\times n}^{\times(l)}\\ E_{l}\\ \mathbb{O}_{q\times n}^{\times(r-l)}\\ \mathbb{O}_{m\times n}^{\times(l)}\\ \sqrt{\frac{1}{r+1}R}\,T(P_{\text{max}})\\ \mathbb{O}_{m\times n}^{\times(r-l)}\end{bmatrix}\in\mathbb{R}^{[(q+m)(r+1)]\times n},\,0\leq l\leq r.$
Since the system $\left[E_{0},E_{1},\dots,E_{r};A,C_{1},\dots,C_{r}\right]$ is stochastically detectable, there exists $\varTheta\in\mathbb{R}^{n\times q}$ such that $\left(A+\varTheta E_{0},C_{1}+\varTheta E_{1},\dots,C_{r}+\varTheta E_{r}\right)$ is mean-square stable. Define
$\hat{\varTheta}=\begin{bmatrix}\varTheta^{\times(r+1)}&-B\sqrt{(r+1)R^{-1}}&-D_{1}\sqrt{(r+1)R^{-1}}&\cdots&-D_{r}\sqrt{(r+1)R^{-1}}\end{bmatrix}\in\mathbb{R}^{n\times[(q+m)(r+1)]},$
then the system $\left(A+BT(P_{\text{max}})+\hat{\varTheta}\hat{E}_{0},C_{1}+D_{1}T(P_{\text{max}})+\hat{\varTheta}\hat{E}_{0},\dots,C_{r}+D_{r}T(P_{\text{max}})+\hat{\varTheta}\hat{E}_{0}\right)$ is mean-square stable. By Lemma 3.5, the system $(A+BT(P_{\text{max}}),\ C_{1}+D_{1}T(P_{\text{max}}),\ \dots,\ C_{r}+D_{r}T(P_{\text{max}}))$ is also mean-square stable. Hence, $P_{\text{max}}$ is a stabilizing solution of ARE (16). By Theorem 5.6.5 in [12], the stabilizing solution is unique, so ARE (16) has a unique stabilizing solution $P\in\overline{\mathbb{S}}_{+}^{n}$ such that $R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\succ 0$. ∎
3.3 Convergence Analysis of Iterative Sequence
Next, we introduce two linear operators associated with the interrelated iterative steps, whose specific properties are analyzed in Lemma 3.8. These favorable properties are crucial for proving the convergence of the algorithm’s iterative process, as they jointly ensure the boundedness of the iterative sequence—a fact explicitly demonstrated in the proof of Theorem 3.9.
For each triple $(P,Z,L)$ satisfying $P\in\mathrm{Dom}\,\mathcal{G}$, $P+Z\in\mathrm{Dom}\,\mathcal{G}$ and $L\in\mathcal{A}$, the linear operators $\mathcal{L}^{*}_{P}$ and $\mathcal{L}^{*}_{P+Z}$ are defined as follows:
$\mathcal{L}^{*}_{P}(Y)=$ $Y\left(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}L\right)+\left(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}L\right)^{\top}Y$ (19)
$+\sum_{l=1}^{r}\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}L\right)^{\top}Y\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}L\right)\,,\forall\,Y\in\mathbb{S}^{n};$
$\mathcal{L}^{*}_{P+Z}(Y)=$ $Y\left(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L\right)+\left(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L\right)^{\top}Y$ (20)
$+\sum_{l=1}^{r}\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P+Z)+D_{l,2}L\right)^{\top}Y\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P+Z)+D_{l,2}L\right)\,,\forall\,Y\in\mathbb{S}^{n},$
where
$\begin{bmatrix}\hat{K}_{1}(P)\\ \hat{K}_{2}(P)\end{bmatrix}=\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}-\begin{bmatrix}K_{1}(0)\\ K_{2}(0)\end{bmatrix},\quad\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}$. (21)
Lemma 3.8.
Assume $0\in\mathrm{Dom}\,\mathcal{G}$, there exists an $L\in\mathcal{A}$ and $P,Z\in\mathbb{S}^{n}$ having the properties:
  • $P,P+Z\in\mathrm{Dom}\,\mathcal{G}$.
  • $P$ and $Z$ satisfy:
    $\mathcal{G}(P)+Z(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}\hat{K}_{2}(P))+(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}\hat{K}_{2}(P))^{\top}Z$ (22)
    $+\sum_{l=1}^{r}(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))$
    $-[B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))]^{\top}R_{22}(P+Z)^{-1}$
    $\quad\times[B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))]=0.$
Let $L\in\mathcal{A}$ and $\tilde{P}_{L}$ be the stabilizing solution of the ARE (7). If $\mathcal{L}^{*}_{P}$ , $\mathcal{L}^{*}_{P+Z}$ are linear operators associated to the triple $(P,Z,L)$ via (19) and (20), then the following assertions hold:
If   $\text{Spec}\,\mathcal{L}^{*}_{P}\subset\mathbb{C}^{-}$, then $\tilde{P}_{L}\succeq P+Z$.
If   $\tilde{P}_{L}\succeq P+Z$, then $\text{Spec}\,\mathcal{L}^{*}_{P+Z}\subset\mathbb{C}^{-}$.
Proof.
First, we prove the Conclusion (i.). From Proposition 3.2, we can obtain:
$\mathcal{G}(P+Z)=$ $\mathcal{L}^{*}_{P}(P+Z)+Q_{L}+\hat{K}^{\top}_{1}(P)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P)+\hat{K}_{1}^{\top}(P)R_{11}\hat{K}_{1}(P)$
$-(J(P)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P)-\hat{K}(P+Z)),$
where $J(P)=\begin{bmatrix}\hat{K}_{1}(P)\\ L\end{bmatrix}$, $\mathcal{L}^{*}_{P}$ is defined in (19), $\hat{K}_{1}(P)$ and $\hat{K}(P)$ is defined in (21). Combining Eq. (22) and Proposition 3.3, we find that $P+Z$ solves the following ARE:
$\mathcal{L}^{*}_{P}(P+Z)+Q_{L}+\hat{K}^{\top}_{1}(P)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P)+\hat{K}_{1}^{\top}(P)R_{11}\hat{K}_{1}(P)$ (23)
$-(J(P)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P)-\hat{K}(P+Z))+(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}$
$\quad\times\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))=0,$
where $\hat{N}_{1}(P,Z)=B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P))+D_{l,2}\hat{K}_{2}(P))$, $\hat{N}_{2}(P,Z)=B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P))+D_{l,2}\hat{K}_{2}(P))$ and $\mathbb{R}_{22}^{\sharp}(P+Z)$ is defined in (12).
Moreover,
$(J(P)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P)-\hat{K}(P+Z))$ (24)
$=(\hat{K}_{1}(P)-\hat{K}_{1}(P+Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)(\hat{K}_{1}(P)-\hat{K}_{1}(P+Z))+H^{\top}_{1}H_{1}=$
$(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))$
$+H_{1}^{\top}H_{1},$
where $H_{1}=\begin{bmatrix}R_{22}(P+Z)^{-\frac{1}{2}}R_{21}(P+Z)&R_{22}(P+Z)^{\frac{1}{2}}\end{bmatrix}(J(P)-\hat{K}(P+Z))$. Next, by virtue of Proposition 3.2, $\tilde{P}_{L}$ can be rewritten as:
$\mathcal{L}^{*}_{P}(\tilde{P}_{L})+Q_{L}+\hat{K}^{\top}_{1}(P)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P)+\hat{K}_{1}^{\top}(P)R_{11}\hat{K}_{1}(P)$ (25)
$-(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L}))=0.$
where $K_{L}(\tilde{P}_{L})=-(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}(B^{\top}_{L}\tilde{P}_{L}+\sum_{l=1}^{r}D^{\top}_{l,1}\tilde{P}_{L}C_{l}+R_{12}L)$.
Subtracting Eq. (23) from Eq. (25) and combining it with Eq. (24) yields:
$\mathcal{L}^{*}_{P}(\tilde{P}_{L}-P-Z)+H^{\top}H,$
$H^{\top}H=H_{1}^{\top}H_{1}-(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L})).$
Since the $\text{Spec}\,\mathcal{L}^{*}_{P}\subset\mathbb{C}^{-}$ and $H^{\top}H\succeq 0$, by the similarly proof of (a.) $\Rightarrow$ (b.) in Lemma 3.5, we have $\tilde{P}_{L}\succeq P+Z$. This completes the proof of Conclusion (i.).
Next, we proceed to prove Conclusion (ii.). Similarly, by Proposition 3.2 and 3.3, and through combining Eq. (13) with Eq. (22), we can derive that $P+Z$ is a solution to the following ARE:
$\mathcal{L}^{*}_{P+Z}(P+Z)+Q_{L}+\hat{K}^{\top}_{1}(P+Z)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P+Z)+\hat{K}_{1}^{\top}(P+Z)R_{11}\hat{K}_{1}(P+Z)$- (26)
$(J(P+Z)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P+Z)-\hat{K}(P+Z))+(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}$
$\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))=0.$
Additionally, by virtue of Proposition 3.2, $\tilde{P}_{L}$ can be rewritten as:
$\mathcal{L}^{*}_{P+Z}(\tilde{P}_{L})+Q_{L}+\hat{K}^{\top}_{1}(P+Z)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P+Z)+\hat{K}_{1}^{\top}(P+Z)R_{11}\hat{K}_{1}(P+Z)$ (27)
$-(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))=0.$
and $(J(P+Z)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P+Z)-\hat{K}(P+Z))=$
$(\hat{K}_{1}(P+Z)-\hat{K}_{1}(P+Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)(\hat{K}_{1}(P+Z)-\hat{K}_{1}(P+Z))+H^{\top}_{2}H_{2}=H^{\top}_{2}H_{2}$ (28)
where $H_{2}=\begin{bmatrix}R_{22}(P+Z)^{-\frac{1}{2}}R_{21}(P+Z)&R_{22}(P+Z)^{\frac{1}{2}}\end{bmatrix}(J(P+Z)-\hat{K}(P+Z))$. Subtracting Eq. (26) from Eq. (28) combining it with Eq. (27) yields:
$\mathcal{L}^{*}_{P+Z}(\tilde{P}_{L}-P-Z)-(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))$
$+H_{2}^{\top}H_{2}-(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}$
$\quad\times(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))=0.$
Let $\Delta=\tilde{P}_{L}-P-Z$, we have: $\mathcal{L}^{*}_{P+Z}\Delta+\sum_{l=0}^{r}\hat{E}_{l}^{\top}\hat{E}_{l}=0,\Delta\succeq 0$, where ($0\leq l\leq r$)
$\hat{E}_{l}=\begin{bmatrix}\mathbb{O}_{m_{1}\times n}^{\times(l)}\\ \sqrt{-\frac{1}{r+1}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})}(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))\\ \mathbb{O}_{m_{1}\times n}^{\times(r-l)}\\ \frac{1}{\sqrt{r+1}}H_{2}\\ \sqrt{-\frac{1}{r+1}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))\end{bmatrix}\in\mathbb{R}^{[m_{2}+m_{1}(r+2)]\times n}.$
Choose
$\hat{\varTheta}=\begin{bmatrix}-\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}}^{\top}B_{1}^{\top}\\ -\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}}^{\top}D_{1,1}^{\top}\\ \dotsb\\ -\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l})^{-1}}^{\top}D_{r,1}^{\top}\\ \mathbb{O}_{n\times m_{2}}^{\top}\\ \mathbb{O}_{n\times m_{1}}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n\times[m_{2}+m_{1}(r+2)]}.$
We obtain that $(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L+\hat{\varTheta}\hat{E}_{0},C_{1,(0)}+D_{1,1}\hat{K}_{1}(P+Z)+D_{1,2}L+\hat{\varTheta}\hat{E}_{1},\dotsb,C_{r,(0)}+D_{r,1}\hat{K}_{1}(P\\ +Z)+D_{r,2}L+\hat{\varTheta}\hat{E}_{r})$ associated with the system $(A_{L}+B_{1}K_{L}(\tilde{P}_{L}),C_{1L}+D_{1,1}K_{L}(\tilde{P}_{L}),\dotsb,C_{rL}+D_{r,1}K_{L}(\tilde{P}_{L}))$ is mean-square stable, which implies the system $\hat{E}_{0},\hat{E}_{1},\dotsb,\hat{E}_{r};A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L,C_{1,(0)}+D_{1,1}\hat{K}_{1}(P+Z)+D_{1,2}L,\dotsb,C_{r,(0)}+D_{r,1}\hat{K}_{1}(P+Z)+D_{r,2}L]$ is stochastically detectable.
By Lemma 3.5, we have $\text{Spec}\,\mathcal{L}^{*}_{P+Z}\subset\mathbb{C}^{-}$. Thus, the proof of Conclusion (ii.) is completed. ∎
Before presenting the next theorem, we first introduce two families of linear operators, $\mathcal{L}^{*}_{P^{(k)}}$ and $\mathcal{L}^{(k,k+1)}$.
$\mathcal{L}^{*}_{P^{(k)}}(Y)=\,Y\left(A_{(0)}+B_{1}\hat{K}_{1}(P^{(k)})+B_{2}L)\right)+\left(A_{(0)}+B_{1}\hat{K}_{1}(P^{(k)})+B_{2}L)\right)^{\top}Y$ (29)
$+\sum_{l=1}^{r}\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P^{(k)})+D_{l,2}L)\right)^{\top}Y\left(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P^{(k)})+D_{l,2}L\right)\,,\quad\forall\,Y\in\mathbb{S}^{n},k=0,1,2,\ldots$
$\mathcal{L}^{(k,k+1)}(Y)=\,Y\left(A_{(k)}+B_{2}T_{(k)}\right)+\left(A_{(k)}+B_{2}T_{(k)}\right)^{\top}Y$ (30)
$+\sum_{l=1}^{r}\left(C_{l,(k)}+D_{l,2}T_{(k)}\right)^{\top}Y\left(C_{l,(k)}+D_{l,2}T_{(k)}\right)\,,\quad\forall\,Y\in\mathbb{S}^{n},k=0,1,2,\ldots$
where $T_{(k)}=-R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}$, $A_{(k)},C_{l,(k)}(1\leq l\leq r),N_{2,(k)}$ is defined in (11) and (12).
Theorem 3.9.
Assume the following conditions hold:
  • $R_{22}\succ 0$ and there exists an $L\in\mathcal{A}$.
  • There exist matrices $E_{l,(0)}(1\leq l\leq r)$ such that: $\sum_{l=0}^{r}E_{l,(0)}^{\top}E_{l,(0)}=Q-S^{\top}(0)R(0)^{-1}S(0)$ and the system
    $\left[E_{0,(0)},E_{1,(0)},\dots,E_{r,(0)};A_{(0)},C_{1,(0)},\ldots,C_{r,(0)}\right]$
    is stochastically detectable.
Then, we have:
The sequences $\{Z^{(k)}\}_{k\geq 0}$, $\{P^{(k)}\}_{k\geq 0}$ are well defined by (8) (9) (10) , and for each $k=0,1,2\ldots$ the following items are fulfilled:
  1. $a_{k}$.
    $\text{Spec}\,\mathcal{L}^{(k,k+1)}\subset\mathbb{C}^{-}$;
  2. $b_{k}$.
    Let $L\in\mathcal{A}$ is arbitrary but fixed and $\tilde{P}_{L}$ be the stabilizing solution to the corresponding ARE (7), then $\tilde{P}_{L}\succeq P^{(k)}+Z^{(k)}$ and $\text{Spec}\,\mathcal{L}_{P^{(k+1)}}\subset\mathbb{C}^{-}$;
  3. $c_{k}$.
    $P^{(k)},P^{(k)}+Z^{(k)}\in\text{Dom }\mathcal{G}$ and $\mathcal{G}(P^{(k)}+Z^{(k)})=-\left(N_{1,(k)}-R_{12}(P^{(k)}+Z^{(k)})R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}\right)^{\top}$
    $\times\mathbb{R}_{22}^{\sharp}(P^{(k)}+Z^{(k)})^{-1}\left(N_{1,(k)}-R_{12}(P^{(k)}+Z^{(k)})R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}\right),$
    where $\mathbb{R}_{22}^{\sharp}(P^{(k)}+Z^{(k)}),N_{1,(k)},N_{2,(k)}$ is defined in (12);
$\lim_{k\rightarrow\infty}P^{(k)}=P^{*}$, where $P^{*}$ is the stabilizing solution of the stochastic GTARE (3).
Proof.
When $k=0$ , $P^{(0)}=0$ and $Z^{(0)}$ is the stabilizing solution to the following ARE:
$ZA_{(0)}+A_{(0)}^{\top}Z+\sum_{l=1}^{r}C_{l,(0)}^{\top}ZC_{l,(0)}+Q-S^{\top}(0)R(0)^{-1}S(0)$ (31)
$-\left(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(0)}\right)^{\top}\left(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,2}\right)^{-1}\left(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(0)}\right)=0.$
To show that $Z^{(0)}$ is well-defined, we first prove the existence unique stabilizing solution to the above equation. Since the set $\mathcal{A}$ is nonempty and set $L^{(0)}=L$, we can find the system $(A_{(0)}+B_{2}L^{(0)},C_{1,(0)}+D_{1,2}L^{(0)},\dots,C_{r,(0)}+D_{r,2}L^{(0)})$ associated with the system $(A_{L},C_{1L},\dots,C_{rL})$ is mean-square stable. This means the system
$A_{(0)},C_{1,(0)},\dotsb,C_{r,(0)};B_{2},D_{1,2},\dotsb,D_{r,2}]$
is mean-square stabilizable. Since the system $E_{0,(0)},E_{1,(0)},\dots,E_{r,(0)};A_{(0)},C_{1,(0)},\dotsb,C_{r,(0)}]$ is stochastically detectable. And the assumption $R_{22}\succ 0$ and $Q-S^{\top}(0)R(0)^{-1}S(0)\succ 0$ ensures that
$0\in\bm{\Gamma}^{[A_{(0)},C_{1,(0)},\dotsb,C_{r,(0)};B_{2},D_{1,2},\dotsb,D_{r,2};\sum_{l=0}^{r}E^{\top}_{l,(0)}E_{l,(0)},R_{22}]}.$
Consequently, by Proposition 3.7, the Eq. (31) admits a unique stabilizing solution $Z^{(0)}$ such that $R_{22}(Z^{(0)})\succ 0$. Thus, $Z^{(0)}$ is well-defined as the unique stabilizing solution to the equation. This further implies that $\text{Spec}\,\mathcal{L}^{(0,1)}\subset\mathbb{C}^{-}$.
Let $L\in\mathcal{A}$ be arbitrary but fixed, $\tilde{P}_{L}$ denote the stabilizing solution to the corresponding ARE (7). Since the operator $\mathcal{L}^{*}_{P^{(0)}}$ is associated with the system $(A_{L},C_{1L}\dotsc,C_{rL})$ is mean-square stable, we have the $\text{Spec}\,\mathcal{L}_{P^{(0)}}\subset\mathbb{C}^{-}$. From (i.) and (ii.) in Lemma 3.8, we get $\tilde{P}_{L}\succeq P^{(0)}+Z^{(0)}$ and $\text{Spec}\,\mathcal{L}_{P^{(1)}}\subset\mathbb{C}^{-}$. Combining $R_{22}(\tilde{P}_{L})\succeq R_{22}\succ 0$ and $R_{11}\preceq R_{11}(\tilde{P}_{L})\prec 0$, we have $P^{(0)}$ and $\tilde{P}_{L}\in\text{Dom }\mathcal{G}$. Since $R_{22}(P^{(0)}+Z^{(0)})\succeq R_{22}\succ 0$ and $R(P^{(0)}+Z^{(0)})\prec R(\tilde{P}_{L})$, we obtain that $R_{11}(P^{(0)}+Z^{(0)})\preceq\mathbb{R}_{22}^{\sharp}(P^{(0)}+Z^{(0)})\preceq\mathbb{R}_{22}^{\sharp}(\tilde{P}_{L})\prec 0$ by using Corollary 4.5 in [16]. It follows that $P^{(0)}+Z^{(0)}\in\text{Dom }\mathcal{G}$. Substituting Eq. (8) into $\mathcal{G}\left(P^{(0)}+Z^{(0)}\right)$ by using Proposition 3.3, then gives:
$\mathcal{G}\left(P^{(0)}+Z^{(0)}\right)=(N_{1,(0)}-R_{12}(Z^{(0)})R_{22}(Z^{(0)})^{-1}N_{2,(0)})^{\top}R^{\sharp}_{22}(Z^{(0)})^{-1}(N_{1,(0)}-R_{12}(Z^{(0)})R_{22}(Z^{(0)})^{-1}N_{2,(0)}).$
This completes the proof of statements $a_{0}-c_{0}$.
Assume k=h1k=h-1, $Z^{(h-1)}$ is well-defined and that $a_{h-1}-c_{h-1}$ hold. We shall prove that $Z^{(h)}$ is well-defined and that $a_{h}-c_{h}$ hold.
For $k=h$ , $Z^{(h)}$ satisfies the following ARE:
$ZA_{(h)}+A_{(h)}^{\top}Z+\sum_{l=1}^{r}C_{l,(h)}^{\top}ZC_{l,(h)}+\sum_{l=0}^{r}E^{\top}_{l,(h)}E_{l,(h)}$ (32)
$-\left(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(h)}\right)^{\top}\left(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,2}\right)^{-1}\left(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(h)}\right)=0,$
where
$E_{l,(h)}=\begin{bmatrix}\mathbb{O}_{m_{1}\times n}^{\times(l)}\\ \sqrt{-\frac{1}{r+1}\mathbb{R}_{22}^{\sharp}\left(P^{(h)}\right)^{-1}}\left(N_{1,(h-1)}-R_{12}(P^{(h)})R_{22}(P^{(h)})^{-1}N_{2,(h-1)}\right)\\ \mathbb{O}_{m_{1}\times n}^{\times(r-l)}\end{bmatrix}\in\mathbb{R}^{[m_{1}(r+1)]\times n},0\leq l\leq r.$
We demonstrate that $Z^{(h)}$ is well-defined. By the Proposition 3.7, proving that $Z^{(h)}$ is well-defined is equivalent to proving that the system $\left[A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)};B_{2},D_{1,2},\dotsb,D_{r,2}\right]$ is mean-square stabilizable and $\left[E_{0,(h)},E_{1,(h)},\dots,E_{r,(h)};A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)}\right]$ is stochastically detectable. To prove the stability of the system $A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)};B_{2},D_{1,2},\dotsb,D_{r,2}]$, it suffices to set $L^{(h)}=L-\hat{K}_{2}(P^{(h)})$, where $\hat{K}_{2}(P)$ is defined in (21). We can then identify the system $\left(A_{(h)}+B_{2}L^{(h)},C_{1,(h)}+D_{1,2}L^{(h)},\dots,C_{r,(h)}+D_{r,2}L^{(h)}\right)$ associated with the operator $\mathcal{L}_{P^{(h)}}$. Since the condition $\text{Spec}\,\mathcal{L}_{P^{(h)}}\subset\mathbb{C}^{-}$, we prove that the system
$\left[A_{(h)},C_{1,(h)},\dots,C_{r,(h)};B_{2},D_{1,2},\dots,D_{r,2}\right]$
is mean-square stabilizable. Let
$\tilde{\varTheta}^{(h)}=\begin{bmatrix}\tilde{\varTheta}_{0}^{(h)}&\dots&\tilde{\varTheta}_{r}^{(h)}\end{bmatrix}\in\mathbb{R}^{n\times[m_{1}(r+1)]},$
where
$\tilde{\varTheta}_{0}^{(h)}=-(B_{1}-B_{2}R_{22}(P^{(h)})^{-1}R_{21}(P^{(h)}))\sqrt{-(r+1)\mathbb{R}_{22}^{\sharp}\left(P^{(h)}\right)^{-1}}$
and
$\tilde{\varTheta}_{l}^{(h)}=-(D_{l1}-D_{l2}R_{22}(P^{(h)})^{-1}R_{21}(P^{(h)}))\sqrt{-(r+1)\mathbb{R}_{22}^{\sharp}\left(P^{(h)}\right)^{-1}},1\leq l\leq r.$
Then we can find the system $(A_{(h)}+\tilde{\varTheta}^{(h)}E_{0,(h)},C_{1,(h)}+\tilde{\varTheta}^{(h)}E_{1,(h)},\dots,C_{r,(h)}+\tilde{\varTheta}^{(h)}E_{r,(h)})$ associated with $\mathcal{L}^{(h-1,h)}$ is mean-square stable. This means the system $E_{0,(h)},E_{1,(h)},\dots,E_{r,(h)};A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)}]$ is stochastically detectable. Similarly, the Eq. (32) admits a unique stabilizing solution $Z^{(l)}$ such that $R_{22}(P^{(l)}+Z^{(l)})\succ 0$ and we also get $\text{Spec}\,\mathcal{L}^{(h,h+1)}\subset\mathbb{C}^{-}$ by using Proposition 3.7.
Since $\text{Spec}\,\mathcal{L}_{P^{(h)}}\subset\mathbb{C}^{-}$, from (i) and (ii) in Lemma 3.8, we get $\tilde{P}_{L}\succeq P^{(h)}+Z^{(h)}$ and $\text{Spec}\,\mathcal{L}_{P^{(h+1)}}\subset\mathbb{C}^{-}$. Similarly, by combining the conditions $R_{22}(P^{(h)}+Z^{(h)})\succ 0$ and $R(P^{(h)}+Z^{(h)})\preceq R(\tilde{P}_{L})$, we apply Corollary 4.5 from [16] to derive $R_{11}(P^{(h)}+Z^{(h)})\preceq\mathbb{R}_{22}^{\sharp}(P^{(h)}+Z^{(h)})\preceq\mathbb{R}_{22}^{\sharp}(\tilde{P}_{L})\prec 0$. Consequently, $P^{(h)}+Z^{(h)}\in\text{Dom }\mathcal{G}$. Substituting (32) into $\mathcal{G}\left(P^{(h)}+Z^{(h)}\right)$ by using Proposition 3.3, then gives:
$\mathcal{G}\left(P^{(h)}+Z^{(h)}\right)=(N_{1,(h)}-R_{12}(P^{(h)}+Z^{(h)})R_{22}(P^{(h)}+Z^{(h)})^{-1}N_{2,(h)})^{\top}R^{\sharp}_{22}(P^{(h)}+Z^{(h)})^{-1}$
$\quad\times(N_{1,(h)}-R_{12}(P^{(h)}+Z^{(h)})R_{22}(P^{(h)}+Z^{(h)})^{-1}N_{2,(h)}).$
Thus, we have proved the statements $a_{h}-c_{h}$. By induction, we conclude that for any $k$, $Z^{(k)}$ is well-defined and $a_{k}-c_{k}$ hold.
In this recursive process, the sequence $\{P^{(k)}\}_{k\geq 0}$ is monotonically non-decreasing and bounded above, so the sequence $\{P^{(k)}\}_{k\geq 0}$ is convergent and $\lim_{k\rightarrow\infty}Z^{(k)}=0$. Set $P^{*}=\lim_{k\rightarrow\infty}P^{(k)}$, given that
$\mathcal{G}(P^{*})=\mathcal{G}(\lim_{k\rightarrow\infty}P^{(k)})=\mathcal{G}\left(\sum_{s=0}^{\infty}Z^{(s)}\right)=\sum_{s=h+1}^{\infty}Z^{(s)}A_{(h)}+\sum_{s=h+1}^{\infty}A_{(h)}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}C_{l,(h)}^{\top}Z^{(s)}C_{l,(h)}$
$-\left(\sum_{s=h+1}^{\infty}B_{2}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(s)}C_{l,(h)}\right)^{\top}\left(R_{22}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(s)}D_{l,2}\right)^{-1}$
$\quad\times\left(\sum_{s=h+1}^{\infty}B_{2}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(s)}C_{l,(h)}\right)+\mathcal{G}\left(\sum_{s=0}^{h}Z^{(s)}\right),\forall h=0,1,2,\cdots.$
Since $\lim_{h\rightarrow\infty}\sum_{s=h+1}^{\infty}Z^{(s)}=0$ and $\lim_{h\rightarrow\infty}\mathcal{G}\left(\sum_{s=1}^{h}Z^{(s)}\right)=$
$\lim_{h\rightarrow\infty}(N_{1,(h)}-R_{12}(P^{(h+1)})R_{22}(P^{(h+1)})^{-1}N_{2,(h)})^{\top}R^{\sharp}_{22}(P^{(h+1)})^{-1}(N_{1,(h)}-R_{12}(P^{(h+1)})R_{22}(P^{(h+1)})^{-1}N_{2,(h)})$
[…]
$0=P(A+B_{2}\tilde{L})+(A+B_{2}\tilde{L})^{\top}P+\sum_{l=1}^{r}(C+D_{l,2}\tilde{L})^{\top}P(C+D_{l,2}\tilde{L})+Q+\tilde{L}^{\top}R_{22}\tilde{L}+\tilde{L}^{\top}S_{2}+S_{2}^{\top}\tilde{L}-$ (33)
$[B^{\top}_{\tilde{W}}P+\sum_{l=1}^{r}D_{l\tilde{W}}^{\top}P(C+D_{l,2}\tilde{L})+S_{\tilde{L}\tilde{W}}]^{\top}(R_{\tilde{W}}+\sum_{l=1}^{r}D_{l\tilde{W}}^{\top}PD_{l\tilde{W}})^{-1}[B^{\top}_{\tilde{W}}P+\sum_{l=1}^{r}D_{l\tilde{W}}^{\top}P(C+D_{l,2}\tilde{L})+S_{\tilde{L}\tilde{W}}$],
[…]
4 Numerical Experiments
To verify the effectiveness of the proposed algorithm, we present a specific numerical example and randomly generate the system parameters as follows:
$A$ $=\begin{bmatrix}0.951718&-0.270753&-0.399002\\ -0.402652&-0.332890&-0.211269\\ 0.865505&0.242714&0.051286\end{bmatrix}$ $Q$ $=\begin{bmatrix}3.030930&-1.047214&-0.749947\\ -1.047214&2.194723&1.675449\\ -0.749947&1.675449&3.333918\end{bmatrix}$
$C_{1}$ $=\begin{bmatrix}0.428394&-0.231063&1.819921\\ 1.636217&-0.730599&1.212696\\ -0.187034&-0.538025&1.193243\end{bmatrix}$ $C_{2}$ $=\begin{bmatrix}0.160803&-1.563720&-1.122872\\ -0.855569&-0.072125&0.710778\\ 0.091879&0.884612&0.977433\end{bmatrix}$
$B_{1}$ $=\begin{bmatrix}0.325576&-0.086447&-0.227595\\ -0.281475&0.347989&-0.397287\\ -0.079578&-0.101469&-0.095411\end{bmatrix}$ $B_{2}$ $=\begin{bmatrix}1.102509&0.029449&0.441306\\ 0.188708&1.055548&0.363029\\ 0.257869&0.036254&1.116683\end{bmatrix}$
$D_{11}$ $=\begin{bmatrix}0.007102&0.008817&0.004927\\ 0.006974&0.000950&0.001090\\ 0.009301&0.004565&0.001537\end{bmatrix}$ $D_{12}$ $=\begin{bmatrix}0.005681&0.002687&0.000159\\ 0.000953&0.009114&0.008537\\ 0.007980&0.009017&0.006756\end{bmatrix}$
$D_{21}$ $=\begin{bmatrix}0.009843&0.001642&0.003074\\ 0.002716&0.001324&0.004221\\ 0.008972&0.003174&0.003310\end{bmatrix}$ $D_{22}$ $=\begin{bmatrix}0.000377&0.005941&0.009387\\ 0.003089&0.006808&0.007371\\ 0.005642&0.006324&0.007449\end{bmatrix}$
$R_{11}$ $=\begin{bmatrix}-1.658649&-0.888904&-0.707647\\ -0.888904&-1.960809&-0.586128\\ -0.707647&-0.586128&-1.471209\end{bmatrix}$ $R_{12}$ $=\begin{bmatrix}0.319075&0.370440&0.509807\\ 0.212195&0.132083&0.857426\\ 0.313025&0.677299&0.000364\end{bmatrix}$
$R_{21}$ $=\begin{bmatrix}0.319075&0.212195&0.313025\\ 0.370440&0.132083&0.677299\\ 0.509807&0.857426&0.000364\end{bmatrix}$ $R_{22}$ $=\begin{bmatrix}1.078023&0.529765&0.411900\\ 0.529765&0.908261&0.312832\\ 0.411900&0.312832&1.135871\end{bmatrix}$
$S_{1}$ $=\begin{bmatrix}0.779145&0.430019&0.484125\\ 0.792295&0.346524&0.258780\\ 0.064140&0.841675&0.116926\end{bmatrix}$ $S_{2}$ $=\begin{bmatrix}0.616144&0.362532&0.803394\\ 0.351730&0.980676&0.676821\\ 0.513458&0.405686&0.268156\end{bmatrix}$
After 13 outer iterations, the algorithm converges to the stabilizing solution of the problem within the specified error tolerance. The norm of the equation residual is $4.9409\times 10^{-6}$. The specific values of the stabilizing solution to GTARE (3) (retained to six decimal places) are presented as follows:
$\begin{bmatrix}4.490400&0.330564&-1.002100\\ 0.330564&4.895330&-0.781543\\ -1.002100&-0.781543&8.616318\end{bmatrix}$.
Figure 1: The figure shows the real parts of three eigenvalues of $Z^{(k-1)}\,(k=1,\dotsb,13)$ evolving with external iterations.
Figure 2: The figure shows the real parts of five eigenvalues of $M_{(k-1)}(k=1,\dotsb,13)$ evolving with external iterations.
Figure 3: This simulation diagram clearly illustrates the dynamic process of a multivariable system under the balanced control strategies of two players.
Figure 1, 2 depicts the evolution of the real parts of eigenvalues across different matrices with external iterations (k=1to 13)(k=1\,to\,13). And the both Figure show a consistent pattern: the real parts of the eigenvalues of $Z^{(k)}$ and $M_{(k)}$ respectively decrease over iterations, approaching 0 by the sixth iteration and continuing to drop, notably featuring a drastic change in the initial $M_{(1)}$ eigenvalues due to one-time disturbances from the opposing player. They collectively reflect an interesting characteristic: When the set $\mathcal{A}$ is allowed to be non-empty, Player 2 starts from a dominant position and injects controls to stabilize the system. Meanwhile, the opposing Player 1 is dissatisfied with the current state. In response, Player 2 makes tentative concessions by continuously granting Player 1 certain advantages. This leads to a gradual reduction in Player 1’s dissatisfaction, which can be intuitively observed from Figure 1,2. Through continuous adjustments, the system approaches its equilibrium position.
Figure 3 shows that the system’s three state variables ($x_{1}$, $x_{2}$, $x_{3}$) undergo a rapid and intense transient process at the initial moment, with the instantaneous pulse of $x_{3}$ being particularly prominent. However, this potential instability trend is quickly suppressed, and all states converge within an extremely short time to reach stability near the zero point. The control input diagrams in the middle and below reveal the reason for this phenomenon: the control commands of the two players exhibit strong, oppositely directed antagonistic pulses in the initial phase. Their interaction does not disrupt the system; instead, it collectively offsets the system’s inherent unstable dynamics, ultimately driving the states toward equilibrium. This is consistent with the conclusions derived from the theoretical research in the Remark 2.6.3 of [28].
Together, these figures verify the effectiveness of our algorithm and indirectly reveal the mechanism by which the algorithm iteratively finds the equilibrium solution.