### v2

# A Unified Computational Approach for Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons

Abstract
This paper proposes a new method for finding closed-loop saddle points in zero-sum linear-quadratic stochastic differential games by decoupling their inherent structure. Specifically, we develop a nested iterative scheme that constructs a monotonically increasing sequence of matrices, thereby decomposing the original problem into interconnected subproblems. By sequentially computing the stabilizing solutions to the algebraic Riccati equations within each subproblem, we obtain the stabilizing solution to the original problem and rigorously establish the convergence of the iterative sequence. A numerical example further validates the effectiveness of the proposed method. To the best of our knowledge, this work extends the classical setting and provides the first general-purpose computational approach for this class of problems.
Keywords: Computational Approach, Stabilizing Solution, Zero-Sum Linear-Quadratic Stochastic Differential Games, Game-Theoretic Algebraic Riccati Equations
1 Introduction
Zero-sum stochastic differential games provide a rigorous mathematical framework for modeling adversarial decision-making in stochastic dynamic systems. Building on the foundational work of [Isaacs1965] and [Pontryagin1967], [Fleming1989] extended the deterministic framework of [Evans1984] to stochastic settings. Within this framework, the zero-sum linear-quadratic stochastic differential games (ZSLQSDG) serve as a canonical simplified model. This model circumvents the significant challenges associated with solving the highly nonlinear Hamilton-Jacobi-Bellman-Isaacs equation, making it a key vehicle for both theoretical and applied research. To date, significant advancements have been achieved in the theoretical study of ZSLQSDG; see, for example, [Mou2006, Sun2014, Yu2015, Sun2016, Sun2021, Zhou2022, Zhang2024, Sun2024].
This paper focuses on the infinite-horizon ZSLQSDG with constant coefficients. For infinite-horizon problems, whether in deterministic or stochastic cases, only the stabilizing solution is of practical interest among all possible solutions. Specifically for this problem, [Sun2016] proves that the existence of closed-loop saddle points is characterized by the solvability of an algebraic Riccati equation (ARE) with a sign-indefinite quadratic term under a certain stabilizing condition. In other words, finding the closed-loop saddle point of the problem requires computing the stabilizing solution to its associated ARE.
ARE featuring a sign-indefinite quadratic term is commonly referred to as Game-Theoretic algebraic Riccati equation (GTARE). For further relevant studies, see [Dragan1997, Rami2000, Dragan2001, Damm2004book, Abou2012book, Dragan2013book, Dragan2020, Zhang2024] and their references. Due to the indefiniteness of its quadratic term, the traditional numerical algorithms for solving conventional ARE are no longer applicable. For deterministic GTARE, [Lanzon2008] proposed a recursive method for the positive stabilizing solution. Unlike deterministic systems, stochastic systems with state-dependent noise complicate stabilizing solution computation. To address this, [Feng2010] developed a method for GTARE arising from a specific ZSLQSDG with state-dependent noise. This approach was later extended by [Dragan2011] to stochastic $H_{\infty}$ problems involving state-dependent and control-dependent noise, and further by [Dragan201501] to stochastic $H_{\infty}$ problems involving state-dependent, control-dependent and disturbance-dependent noise. [Ivanov2018] investigated a class of GTAREs arising in stochastic $H_{\infty}$ problems with state-dependent noise. The study explored three effective approaches for computing the stabilizing solutions of these equations and performed numerical comparisons of the iterative methods. Further relevant computational studies can be found in [Dragan2008, Dragan201502, Ivanov2015, Dragan2017, Aberkane2023, Sun2023] and their references.
To the best of our knowledge, no computational method has been reported in the literature for ZSLQSDG with state-dependent and both-control-dependent noise. This paper proposes a new method for ZSLQSDG by decoupling its inherent structure. Specifically, we construct a monotonically increasing sequence of matrices by extending the defect correction method proposed in [Mehrmann1988], thereby decomposing the ZSLQSDG problem into a series of interconnected subproblems. These subproblems are solved sequentially by identifying the stabilizing solutions to their associated AREs, ultimately yielding the stabilizing solutions for the original problem. The main contributions of this work are summarized as follows:
  • First General Computational Approach: Previous research has only tackled special cases of this problem, this work overcomes the inherent coupling challenges and provides the first general computational method for finding the closed-loop saddle points of ZSLQSDG.
  • Extended Applicability: The computational method extends the classical setting and is applicable to three categories of GTAREs: deterministic GTAREs, stochastic $H_{\infty}$-type GTAREs, and more general stochastic GTAREs arising from ZSLQSDG.
  • Theoretical Guarantees: This study establishes rigorous theoretical guarantees for the nested iterative scheme, including the existence, uniqueness and boundedness of the stabilizing solutions to AREs with sign-definite quadratic terms, along with a comprehensive convergence analysis.
The remainder of this paper is organized as follows. Section 2 covers the mathematical framework of ZSLQSDG, a decoupled representation method, and the iteration process of the computational method. Section 3 presents the main results, including the key structural properties of the problem, the criteria for the existence and uniqueness of the stabilizing solutions to the subproblems, and the proofs for the boundedness and convergence of the iterative sequence. Section 4 details a numerical example.
2 Preliminary
2.1 Notation
Let us first introduce the following notation:
  • $\mathbb{R}^{n}$: $n$-dimensional real Euclidean space; $\mathbb{C}^{-}$: the set of complex numbers with negative real part; $\mathbb{R}^{n\times m}$: the space of $n\times m$ real matrices; $\mathbb{S}^{n}$: the set of all $n\times n$ symmetric matrices; $\overline{\mathbb{S}}_{+}^{n}$: the set of all $n\times n$ symmetric positive semi-definite matrices; $\mathbb{S}_{+}^{n}$: the set of all $n\times n$ symmetric positive definite matrices.
  • $I_{n}$: the identity matrix of size $n$; $\mathbb{O}_{n\times m}$: the null matrix of size $n\times m$. It can be simplified as 0 when no ambiguity is generated; $A_{m\times n}^{\times(l)}$ denotes the object represented by $A_{m\times n}$ arranged repeatedly for $l$ times.
  • $A^{\top}$: the transpose of the matrix $A$; $A^{\dagger}$: the Moore-Penrose pseudoinverse of the matrix $A$; $\langle\cdot,\cdot\rangle$: the inner product on a Hilbert space. If $A\in\mathbb{S}_{+}^{n}$ (resp., $A\in\overline{\mathbb{S}}_{+}^{n}$), we write $A\succ 0$ (resp., $A\succeq 0$). For any $A,B\in\mathbb{S}^{n}$, we use the notation $A\succ B$ (resp., $A\succeq B$) to indicate that $A-B\succ 0$ (resp., $A-B\succeq 0$).
[…]
2.2 Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a complete filtered probability space on which a r-dimensional standard Brownian motion $W=\{W(t)^{\top}=(w_{1}(t),\dotsb,w_{r}(t));t\geq 0\}$ is defined with $\mathbb{F}=\{\mathcal{F}_{t}\}_{t\geq 0}$ being the usual augmentation of the natural filtration generated generated by $W$. We consider the following controlled linear stochastic differential equation (SDE) on $0,\infty)$:
$\begin{cases}&dX(t)=[AX(t)+B_{1}u_{1}(t)+B_{2}u_{2}(t)]dt\\ &\quad\quad\quad+\sum_{l=1}^{r}[C_{l}X(t)+D_{l,1}u_{1}(t)+D_{l,2}u_{2}(t)]dw_{l}(t)\\ &X(0)=x\end{cases},$ (1)
in which $x\in\mathbb{R}^{n}$, $A,C_{l}\in\mathbb{R}^{n\times n}$ and $B_{i},D_{l,i}\in\mathbb{R}^{n\times m_{i}}$ ($i=1,2;1\leq l\leq r$). In the above, the state process $X$ is an $n$-dimensional vector, player 1 $u_{1}$ and player 2 $u_{2}$ are an $m_{1}$-dimensional vector and an $m_{2}$-dimensional vector, respectively.
In this paper, Player 1 and Player 2 share the same performance functional:
$J(x;u_{1},u_{2})=\mathbb{E}\int_{0}^{\infty}\Bigg[\Bigg\langle\begin{pmatrix}Q&S_{1}^{\top}&S_{2}^{\top}\\ S_{1}&R_{11}&R_{12}\\ S_{2}&R_{21}&R_{22}\end{pmatrix}\begin{pmatrix}X(t)\\ u_{1}(t)\\ u_{2}(t)\end{pmatrix},\begin{pmatrix}X(t)\\ u_{1}(t)\\ u_{2}(t)\end{pmatrix}\Bigg\rangle\Bigg]dt,$ (2)
the weighting coefficients in (2) satisfy:
$Q\in\mathbb{S}^{n},\quad R_{21}^{\top}=R_{12}\in\mathbb{R}^{m_{1}\times m_{2}},\quad S_{i}\in\mathbb{R}^{m_{i}\times n},R_{ii}\in\mathbb{S}^{m_{i}}(i=1,2).$
For $(x,u_{1},u_{2})\in\mathbb{R}^{n}\times\mathcal{U}_{1}\times\mathcal{U}_{2}$, the solution $X(\cdot;x,u_{1},u_{2})$ to the SDE in (1) may only exist in $\mathcal{X}_{loc}[0,\infty)$, which renders $J(x;u_{1},u_{2})$ ill-defined. We define the set of admissible controls as $\mathcal{U}_{ad}(x)=\{(u_{1},u_{2})\in\mathcal{U}_{1}\times\mathcal{U}_{2}\mid X(\cdot;x,u_{1},u_{2})\in\mathcal{X}[0,\infty)\}$. A pair $(u_{1},u_{2})\in\mathcal{U}_{ad}(x)$ is called an admissible control pair for the initial state $x$, and the corresponding $X(\cdot;x,u_{1},u_{2})$ is referred to as the admissible state process. In this case, $J(x,u_{1},u_{2})$ is clearly well-defined.
In this zero-sum game, Player 1 (the maximizer) selects control $u_{1}$ to maximize (2), while Player 2 (the minimizer) chooses $u_{2}$ to minimize the same function. The problem is to find an admissible control pair $(u_{1}^{*},u_{2}^{*})$ that both players can accept. We denote the above-mentioned problem as $({SDG})^{\,0}_{\infty}$ for short. For a description of the $({SDG})^{\,0}_{\infty}$ problem, refer to [Sun2020_book] . More detailed information can be found therein.
The $({SDG})^{\,0}_{\infty}$ problem corresponds to following stochastic GTARE:
$\begin{cases}Q(P)-S(P)^{\top}R(P)^{\dagger}S(P)=0\\ \mathcal{R}(S(P))\subseteq\mathcal{R}(R(P))\\ R_{11}(P)\preceq 0,\quad R_{22}(P)\succeq 0\end{cases},$ (3)
where $\mathcal{R}(M)$ denotes the range of a matrix $M$,
$Q(P)=PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q,$ (4)
$S(P)=\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}=\begin{bmatrix}B_{1}^{\top}P+\sum_{l=1}^{r}D_{l,1}^{\top}PC_{l}+S_{1}\\ B_{2}^{\top}P+\sum_{l=1}^{r}D_{l,2}^{\top}PC_{l}+S_{2}\\ \end{bmatrix}$,
$R(P)=\begin{bmatrix}R_{11}(P)&R_{12}(P)\\ R_{21}(P)&R_{22}(P)\\ \end{bmatrix},\quad R_{ij}(P)=R_{ij}+\sum_{l=1}^{r}D_{l,i}^{\top}PD_{l,j}\,(i,j=1,2).$
Remark 2.1.
For the stochastic GTARE (3), the general theoretical framework of $({SDG})^{\,0}_{\infty}$ only requires that $R(P)$ satisfies $\mathcal{R}(S(P))\subseteq\mathcal{R}(R(P))$ in [Sun2020_book]. However, in our subsequent discussion, we impose the slightly stronger condition that $R(P)$ be of full rank. This requirement is well-motivated as it allows us to focus on the core structural properties of the $({SDG})^{\,0}_{\infty}$. We therefore restrict our analysis to the case where $R(P)$ is full rank.
Definition 2.2.
The system
$\begin{cases}dX(t)=AX(t)dt+\sum_{l=1}^{r}C_{l}X(t)dw_{l}(t)\\ X(0)=x\end{cases},$
donated as $(A,C_{1},\dotsb,C_{r})$ is called mean-square stable such that the solution satisfies
$\lim_{t\to\infty}\mathbb{E}[X(t)^{\top}X(t)]=0\,\,\text{for every initial state}\,\,x\in\mathbb{R}^{n}.$
A concept related to the above admissible control pair is that of mean-square stabilizers. We define the set of all such stabilizers associated with the system (1) as
$\mathcal{K}=\left\{\begin{aligned} &\text{All of}\,\,\,(\varTheta_{1},\varTheta_{2})\in\mathbb{R}^{m_{1}\times n}\times\mathbb{R}^{m_{2}\times n}\,\text{such that the system}\,(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2},\\ &C_{1}+D_{1,1}\varTheta_{1}+D_{1,2}\varTheta_{2},\dotsb,C_{r}+D_{r,1}\varTheta_{1}+D_{r,2}\varTheta_{2})\text{ is mean-square stable}.\end{aligned}\right\}.$
Let $(\varTheta_{1},\varTheta_{2})\in\mathcal{K}$, the feedback control $U(\cdot)=\begin{bmatrix}\varTheta_{1}X(\cdot)\\ \varTheta_{2}X(\cdot)\end{bmatrix}$ is called mean-square stabilizing. Moreover, for $\varTheta_{i}\in\mathbb{R}^{m_{i}\times n}(i=1,2)$, we let
$\mathcal{K}_{1}(\varTheta_{2})=\left\{\varTheta_{1}\in\mathbb{R}^{m_{1}\times n}:(\varTheta_{1},\varTheta_{2})\in\mathcal{K}\right\},$
$\mathcal{K}_{2}(\varTheta_{1})=\left\{\varTheta_{2}\in\mathbb{R}^{m_{2}\times n}:(\varTheta_{1},\varTheta_{2})\in\mathcal{K}\right\}.$
We refer to [Rami2000, Sun2020_book] for the above definition.
Definition 2.3 ([Sun2020_book]).
A 2-tuple pair $(\varTheta_{1}^{*},\varTheta_{2}^{*})\in\mathbb{R}^{m_{1}\times n}\times\mathbb{R}^{m_{2}\times n}$ is called a closed-loop saddle point of Problem $({SDG})^{\,0}_{\infty}$ if $(\varTheta_{1}^{*},\varTheta_{2}^{*})\in\mathcal{K}$ and
$J(x,\varTheta_{1}X(\cdot),\varTheta_{2}^{*}X(\cdot))\leq J(x,\varTheta_{1}^{*}X(\cdot),\varTheta_{2}^{*}X(\cdot))\leq J(x,\varTheta_{1}^{*}X(\cdot),\varTheta_{2}X(\cdot)),$
for all $x\in\mathbb{R}^{n},(\varTheta_{1},\varTheta_{2})\in\mathcal{K}_{1}(\varTheta_{2}^{*})\times\mathcal{K}_{2}(\varTheta_{1}^{*})$.
Definition 2.4 ([Sun2020_book]).
A matrix $P\in\mathbb{S}^{n}$ is called a stabilizing solution of stochastic GTARE (3) if $P$ satisfies (3) and $(K_{1}(P),K_{2}(P))\in\mathcal{K}$, where $\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{\top}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}$ and $\mathcal{K}$ is the set of all such stabilizers associated with the system (1).
Remark 2.5.
Theorem 2.6.7 in [Sun2020_book] directly establishes the relationship between $({SDG})^{\,0}_{\infty}$ and the stochastic GTARE. Based on this result, we transform the problem of finding the closed-loop saddle point of $({SDG})^{\,0}_{\infty}$ into the equivalent problem of finding the stabilizing solution $P$ to the stochastic GTARE (3).
To facilitate subsequent analysis of the problem, we introduce the following decoupling representation method. Setting
$\begin{bmatrix}v_{1}(t)\\ v_{2}(t)\end{bmatrix}=\begin{bmatrix}u_{1}(t)\\ u_{2}(t)\end{bmatrix}-\begin{bmatrix}K_{1}(0)\\ K_{2}(0)\end{bmatrix},\quad\begin{bmatrix}K_{1}(0)\\ K_{2}(0)\end{bmatrix}=-R(0)^{-1}\begin{bmatrix}S_{1}(0)\\ S_{2}(0)\end{bmatrix}$,
and $v_{2}(t)=Lx(t)$ in $({SDG})^{\,0}_{\infty}$, we obtain
$\begin{cases}dX(t)=[A_{L}X(t)+B_{1}v_{1}(t)]dt+\sum_{l=1}^{r}[C_{lL}X(t)+D_{l,1}v_{1}(t)]dw_{l}(t)\\ x(0)=x\end{cases}$
in which $x\in\mathbb{R}^{n}$, and
$J_{L}(x;v_{1})=\mathbb{E}\int_{0}^{\infty}\Bigg[\Bigg\langle\begin{pmatrix}Q_{L}&S^{\top}_{L}\\ S_{L}&R_{11}\end{pmatrix}\begin{pmatrix}X(t)\\ v_{1}(t)\end{pmatrix},\begin{pmatrix}X(t)\\ v_{1}(t)\end{pmatrix}\Bigg\rangle\Bigg]dt,$
where
$\begin{cases}&A_{L}=A+B_{1}K_{1}(0)+B_{2}K_{2}(0)+B_{2}L\\ &C_{lL}=C_{l}+D_{l,1}K_{1}(0)+D_{l,2}K_{2}(0)+D_{l,2}L,1\leq l\leq r\\ &Q_{L}=Q-S^{\top}(0)R(0)^{-1}S(0)+L^{\top}R_{22}L,\quad S_{L}=R_{12}L\end{cases}.$
The corresponding ARE is
$PA_{L}+A_{L}^{\top}P+\sum_{l=1}^{r}C_{lL}^{\top}PC_{lL}+Q_{L}-(B_{1}^{\top}P+\sum_{l=1}^{r}D_{l,1}^{\top}PC_{lL}+S_{L})^{\top}$ (5)
$\quad\times(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}PD_{l,1})^{-1}(B_{1}^{\top}P+\sum_{l=1}^{r}D_{l,1}^{\top}PC_{lL}+S_{L})=0.$
Throughout this work, $\mathcal{A}$ stands for the set of $L$, where $L\in\mathbb{R}^{m_{2}\times n}$ satisfy:
  • The system $(A_{L},C_{1L},\dotsb,C_{rL})$ is mean-square stable.
  • The corresponding ARE (5) has a stabilizing solution $\tilde{P}_{L}$, satisfying the sign conditions
    $R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1}\prec 0.$
2.3 Nested Iterative Scheme
In this paper, we aim to solve for the closed-loop saddle points numerically for the $({SDG})^{\,0}_{\infty}$ problem with given parameter. Consider the sequences $\{P^{(k)}\}_{k\geq 0}$ and $\{Z^{(k)}\}_{k\geq 0}$ constructed according to the following procedure:
1. For $k=0$, set $P^{(0)}=0$, and let $Z^{(0)}$ be the unique stabilizing solution to the following ARE with sign-definite quadratic term:
$Z^{(0)}A_{(0)}+A_{(0)}^{\top}Z^{(0)}+\sum_{l=1}^{r}C_{l,(0)}^{\top}Z^{(0)}C_{l,(0)}+M_{(0)}-(B_{2}^{\top}Z^{(0)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}C_{l,(0)})^{\top}$ (6)
$\quad\times(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}D_{l,2})^{-1}(B_{2}^{\top}Z^{(0)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(0)}C_{l,(0)})=0,$
where $A_{(0)}=A-\begin{bmatrix}B_{1}&B_{2}\end{bmatrix}R(0)^{-1}S(0),\,\,C_{l,(0)}=C_{l}-\begin{bmatrix}D_{l,1}&D_{l,2}\end{bmatrix}R(0)^{-1}S(0)(1\leq l\leq r)$,   $M_{(0)}=Q-S^{\top}(0)R(0)^{-1}S(0)$.
2. For $k\geq 1$, update $P^{(k)}$ via:
$P^{(k)}=P^{(k-1)}+Z^{(k-1)},$ (7)
and solving $Z^{(k)}$, which is the unique stabilizing solution to the following ARE with sign-definite quadratic term:
$Z^{(k)}A_{(k)}+A_{(k)}^{\top}Z^{(k)}+\sum_{l=1}^{r}C_{l,(k)}^{\top}Z^{(k)}C_{l,(k)}+M_{(k)}-(B_{2}^{\top}Z^{(k)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}C_{l,(k)})^{\top}$ (8)
$\quad\times(R_{22}(P^{(k)})+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}D_{l,2})^{-1}(B_{2}^{\top}Z^{(k)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k)}C_{l,(k)})=0,$
the matrices $A_{(k)},C_{l,(k)}$ and $M_{(k)}$ evolve as:
$\begin{cases}A_{(k)}=A_{(k-1)}-\begin{bmatrix}B_{1}&B_{2}\end{bmatrix}R(P^{(k)})^{-1}\begin{bmatrix}N_{1,(k-1)}\\ N_{2,(k-1)}\end{bmatrix}\\ C_{l,(k)}=C_{l,(k-1)}-\begin{bmatrix}D_{l1}&D_{l2}\end{bmatrix}R(P^{(k)})^{-1}\begin{bmatrix}N_{1,(k-1)}\\ N_{2,(k-1)}\end{bmatrix},\quad 1\leq l\leq r\\ M_{(k)}=[N_{1,(k-1)}-R_{12}(P^{(k)})R_{22}(P^{(k)})^{-1}N_{2,(k-1)}]^{\top}R^{\sharp}_{22}(P^{(k)})^{-1}\\ \quad\quad\quad\times[N_{1,(k-1)}-R_{12}(P^{(k)})R_{22}(P^{(k)})^{-1}N_{2,(k-1)}]\end{cases}$ (9)
where $R^{\sharp}_{22}(P)=R_{11}(P)-R_{12}(P)R_{22}(P)^{-1}R_{21}(P),$ (10)
and $\begin{bmatrix}N_{1,(k-1)}\\ N_{2,(k-1)}\end{bmatrix}=\begin{bmatrix}B_{1}^{\top}Z^{(k-1)}+\sum_{l=1}^{r}D_{l,1}^{\top}Z^{(k-1)}C_{l,(k-1)}\\ B_{2}^{\top}Z^{(k-1)}+\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(k-1)}C_{l,(k-1)}\end{bmatrix}$.
In the subsequent content, we will show:
  • The sequences $\{P^{(k)}\}_{k\geq 0}$ and $\{Z^{(k)}\}_{k\geq 0}$ are well-defined. The AREs in (6) and (8) admits unique stabilizing solutions $Z^{(k)}\succeq 0(k=0,1,2,\dotsb)$ such that $\tilde{P}_{L}\succeq\sum^{h}_{k=0}Z^{(k)}(h=0,1,2,\dotsb)$, where $\tilde{P}_{L}$ be the stabilizing solution of the corresponding ARE (5).
  • These sequences are convergent and we have $\lim_{k\to\infty}P^{(k)}=\tilde{P}$, where $\tilde{P}$ is the stabilizing solution to the stochastic GTARE (3).
3 The Main Results
3.1 Analysis of Structural Characteristics for $({SDG})^{\,0}_{\infty}$
From stochastic GTARE (3), we can define a mapping $\mathcal{G}:\mathrm{Dom}\,\mathcal{G}\to\mathbb{S}^{n}$ that satisfies the following relation:
$\mathcal{G}(P)=PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q-S(P)^{\top}R(P)^{-1}S(P).$ (11)
The nonlinear function $\mathcal{G}$ is not defined over the entire space $\mathbb{S}^{n}$. Instead, it is well-defined on the subset
$\mathrm{Dom}\,\mathcal{G}=\left\{P\in\mathbb{S}^{n}|R_{22}(P)\succ 0\,\,\text{and}\,\,R_{11}(P)\prec 0\right\}.$
The aforementioned mapping $\mathcal{G}$ serves as the fundamental foundation for constructing the nested iterative scheme. We begin by detailing its general properties, which collectively ensure the computational method’s ability to operate iteratively.
First, we define two operators that will be used in the subsequent content:
$\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix},\begin{bmatrix}N_{1}(P,Z)\\ N_{2}(P,Z)\end{bmatrix}=\begin{bmatrix}B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\\ B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z\bigl(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P)\bigr)\end{bmatrix}$, (12)
where $B_{i},D_{l,i}(i=1,2;1\leq l\leq r)$ is defined in (1), and $R(P),S_{1}(P),S_{2}(P)$ is defined in (4).
Proposition 3.1.
Let $P$ and $Z$ be such that $P,P+Z\in\mathrm{Dom}\,\mathcal{G}$. Then the feedback gains satisfy the following relation:
$\begin{bmatrix}K_{1}(P+Z)\\ K_{2}(P+Z)\end{bmatrix}=\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}-R(P+Z)^{-1}\begin{bmatrix}N_{1}(P,Z)\\ N_{2}(P,Z)\end{bmatrix}$.
Proof.
Define the increment $\Delta=\begin{bmatrix}K_{1}(P+Z)\\ K_{2}(P+Z)\end{bmatrix}-\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}$. By leveraging the definitions of $K_{1}(\cdot),K_{2}(\cdot)$ in (12), and multiplying both sides of the equation by $R(P+Z)$, we obtain:
$R(P+Z)\Delta=R(P+Z)R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}-\begin{bmatrix}S_{1}(P+Z)\\ S_{2}(P+Z)\end{bmatrix}$.
Note that,
$R(P+Z)=R(P)+\Delta R,\Delta R=\begin{bmatrix}\sum_{l=1}^{r}D_{l,1}^{\top}ZD_{l,1}&\sum_{l=1}^{r}D_{l,1}^{\top}ZD_{l,2}\\ \sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,1}&\sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,2}\end{bmatrix}$.
Using the identity, $R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}=-\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}$, we have
$R(P+Z)\Delta=\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}+\Delta R\cdot R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}-\begin{bmatrix}S_{1}(P+Z)\\ S_{2}(P+Z)\end{bmatrix}=-\begin{bmatrix}N_{1}(P,Z)\\ N_{2}(P,Z)\end{bmatrix}$.
Multiplying both sides by $R(P+Z)^{-1}$ yields the relation. ∎
Proposition 3.2 ([Dragan2013book]).
For any $P\in\mathrm{Dom}\,\mathcal{G}$ and for all $\varTheta_{1}\in\mathbb{R}^{m_{1}\times n},\varTheta_{2}\in\mathbb{R}^{m_{2}\times n}$, we have
$\mathcal{G}(P)=P(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2})+(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2})^{\top}P+Q$
$+\sum_{l=1}^{r}(C_{l}+D_{l,1}\varTheta_{1}+D_{l,2}\varTheta_{2})^{\top}P(C_{l}+D_{l,1}\varTheta_{1}+D_{l,2}\varTheta_{2})+\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}^{\top}R(0)\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}$
$+\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}^{\top}\begin{bmatrix}S_{1}\\ S_{2}\end{bmatrix}+\begin{bmatrix}S_{1}\\ S_{2}\end{bmatrix}^{\top}\begin{bmatrix}\varTheta_{1}\\ \varTheta_{2}\end{bmatrix}-\begin{bmatrix}K_{1}(P)-\varTheta_{1}\\ K_{2}(P)-\varTheta_{2}\end{bmatrix}^{\top}R(P)\begin{bmatrix}K_{1}(P)-\varTheta_{1}\\ K_{2}(P)-\varTheta_{2}\end{bmatrix}$.
Proposition 3.3.
Let $P$, $Z$ satisfy $P,P+Z\in\mathrm{Dom}\,\mathcal{G}$. Then the following identity holds:
$\mathcal{G}(P+Z)=\mathcal{G}(P)+Z(A+B_{1}K_{1}(P)+B_{2}K_{2}(P))+(A+B_{1}K_{1}(P)+B_{2}K_{2}(P))^{\top}Z$
$+\sum_{l=1}^{r}(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))^{\top}Z(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))$
$-N(P,Z)^{\top}R(P+Z)^{-1}N(P,Z),\quad\text{where}\,\,N(P,Z)=\begin{bmatrix}N_{1}(P,Z)\\ N_{2}(P,Z)\end{bmatrix}$.
Proof.
Starting from the definition of $\mathcal{G}$, $\mathcal{G}(P+Z)=(P+Z)A+A^{\top}(P+Z)+\sum_{l=1}^{r}C_{l}^{\top}(P+Z)C_{l}+Q-S(P+Z)^{\top}R(P+Z)^{-1}S(P+Z).$
From Proposition 3.2, expanding and rearranging terms, we obtain $\mathcal{G}(P+Z)=$
$PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q+K(P)^{\top}R(P)K(P)+S(P)^{\top}K(P)+K(P)^{\top}S(P)$
$+Z(A+B_{1}K_{1}(P)+B_{2}K_{2}(P))+(A+B_{1}K_{1}(P)+B_{2}K_{2}(P))^{\top}Z$
$+\sum_{l=1}^{r}(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))^{\top}Z(C_{l}+D_{l,1}K_{1}(P)+D_{l,2}K_{2}(P))$
$-(K(P+Z)-K(P))^{\top}R(P+Z)(K(P+Z)-K(P)),\,\,\text{where}\,\,K(P)=\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}$.
From Proposition 3.1, we derive:
$(K(P+Z)-K(P))^{\top}R(P+Z)(K(P+Z)-K(P))=N(P,Z)^{\top}R(P+Z)^{-1}N(P,Z).$
Substitute the above term into the rearranged expression of $\mathcal{G}(P+Z)$, and the target identity is thus verified to hold. ∎
3.2 Existence and Uniqueness of Stabilizing Solution for ARE with Sign-Definite Quadratic Term
Based on the structure of $\mathcal{G}$ and Proposition 3.3, a series of interconnected sub-algebraic Riccati equations with sign-definite quadratic terms can be constructed, as detailed in Section 2.3. In each iteration step, the existence and uniqueness of stabilizing solutions to these equations are of particular importance. In what follows, we establish a criterion for such equations and separately present some related properties—both of which will be repeatedly employed in proving the iteration sequence’s convergence.
Definition 3.4 ([Dragan2013book]).
Given the following stochastic observation system:
$\begin{cases}dX(t)=AX(t)dt+\sum_{l=1}^{r}C_{l}X(t)dw_{l}(t)\\ dY(t)=E_{0}X(t)dt+\sum_{l=1}^{r}E_{l}X(t)dw_{l}(t)\end{cases},$
where $E_{l}\in R^{q\times n}(0\leq l\leq r)$, we denote this system by $\left[E_{0},E_{1},\dotsb,E_{r};A,C_{1},\dotsb,C_{r}\right]$. It is said to be stochastically detectable if there exists a constant matrix $\varTheta\in\mathbb{R}^{n\times q}$ such that the system $(A+\varTheta E_{0},\ C_{1}+\varTheta E_{1},\dotsb,C_{r}+\varTheta E_{r})$ is mean-square stable.
Lemma 3.5.
If the system $\left[E_{0},E_{1},\dotsb,E_{r};A,C_{1},\dotsb,C_{r}\right]$ is stochastically detectable, then the following statements are equivalent:
  • (a).
    The system $(A,C_{1},\dotsb,C_{r})$ is mean-square stable. Let $\mathcal{L}^{*}$ denote the linear operator associated with this system, defined by $\mathcal{L}^{*}(P)=PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l};\,\forall P\in\mathbb{S}^{n}$. Then all eigenvalues of $\mathcal{L}^{*}$ lie in the left half-plane, i.e., $\operatorname{Spec}\mathcal{L}^{*}\subset\mathbb{C}^{-}$.
  • (b).
    There exists a matrix $P\in\overline{\mathbb{S}}_{+}^{n}$ satisfying the matrix equation $PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+\sum_{l=0}^{r}E_{l}^{\top}E_{l}=0$.
Proof.
(a) $\Rightarrow$ (b): Since $\sum_{l=0}^{r}E_{l}^{\top}E_{l}\succeq 0$ and the system $(A,C_{1},\dotsb,C_{r})$ is mean-square stable, it follows from Theorem 3.2.2. and Theorem 2.7.5 in [Dragan2013book] that the equation $PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+\sum_{l=0}^{r}E_{l}^{\top}E_{l}=0$ has a solution $P\in\overline{\mathbb{S}}_{+}^{n}$.
(b) $\Rightarrow$ (a): This conclusion follows directly from Remark 4.1.5 in [Dragan2013book]. For a detailed argument, see Theorem 4.1.7 and Remark 4.1.5 in the same reference. Furthermore, if the system $(A,C_{1},\dotsb,C_{r})$ is mean-square stable, then $\operatorname{Spec}\mathcal{L}^{*}\subset\mathbb{C}^{-}$. This is a consequence of Theorem 3.2.2. and Theorem 2.7.7 in [Dragan2013book]. ∎
Now we consider ARE of the form
$PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q-(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l})^{\top}(R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l})^{-1}(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l})=0,$ (13)
where $B,D_{l}\in\mathbb{R}^{n\times m}$ for $1\leq l\leq r$, and $R\in\mathbb{R}^{m\times m}$. The remaining parameters follow the previous definitions.
Define the operator $\bm{\Lambda}^{[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r};Q,R]}:\mathbb{S}^{n}\rightarrow\mathbb{S}^{n+m}$ associated with the ARE (13) as
$\bm{\Lambda}^{[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r};Q,R]}(P)=\begin{bmatrix}PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q&PB+\sum_{l=1}^{r}C_{l}^{\top}PD_{l}\\ B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}&R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\end{bmatrix}$.
Define the set $\bm{\Gamma}^{[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r};Q,R]}$ related to ARE (13) as
$\left\{P\in\mathbb{S}^{n}\,\middle|\,\bm{\Lambda}^{[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r};Q,R]}(P)\succeq 0,\ R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\succ 0\right\}.$
Definition 3.6 ([Dragan2013book]).
The system associated with the ARE (13) denoted as $A,C_{1},\dotsb,C_{r};\,B,D_{1},\\ \dotsb,D_{r}]$ is said to be mean-square stabilizable if there exists a constant matrix $\varTheta\in\mathbb{R}^{n\times m}$ such that the system $\left(A+B\varTheta,C_{1}+D_{1}\varTheta,\dotsb,C_{r}+D_{r}\varTheta\right)$ is mean-square stable.
Proposition 3.7.
Suppose the parameters of the ARE (13) satisfy the following:
  • $0\in\bm{\Gamma}^{[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r};Q,R]}$.
  • The system $A,C_{1},\dotsb,C_{r};\ B,D_{1},\dotsb,D_{r}]$ is mean-square stabilizable.
  • There exists a matrix set $\{E_{0},E_{1},\dotsb,E_{r}\}$ satisfying $\sum_{l=0}^{r}E_{l}^{\top}E_{l}=Q$ such that the system
    $E_{0},E_{1},\dotsb,E_{r};A,C_{1},\dotsb,C_{r}]$
    is stochastically detectable.
Then the ARE (13) admits a unique stabilizing solution $P\in\overline{\mathbb{S}}_{+}^{n}$ such that $R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\succ 0$.
Proof.
By Theorem 4.7 in [Drgan2004], along with the condition $0\in\bm{\Gamma}$ and the mean-square stabilizability of $\left[A,C_{1},\dotsb,C_{r};B,D_{1},\dotsb,D_{r}\right]$, the ARE (13) has a maximal solution $P_{\text{max}}\in\overline{\mathbb{S}}_{+}^{n}$ such that $R+\sum_{l=1}^{r}D_{l}^{\top}P_{\text{max}}D_{l}\succ 0$.
Using Proposition 3.2, the ARE for $P_{\text{max}}$ can be rewritten as
$P_{\text{max}}\left(A+BT(P_{\text{max}})\right)+\left(A+BT(P_{\text{max}})\right)^{\top}P_{\text{max}}$
$+\sum_{l=1}^{r}\left(C_{l}+D_{l}T(P_{\text{max}})\right)^{\top}P_{\text{max}}\left(C_{l}+D_{l}T(P_{\text{max}})\right)+\sum_{l=0}^{r}\hat{E}_{l}^{\top}\hat{E}_{l}=0,$
where $T(P)=-\left(R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\right)^{-1}\left(B^{\top}P+\sum_{l=1}^{r}D_{l}^{\top}PC_{l}\right)$, and
$\hat{E}_{l}=\begin{bmatrix}\mathbb{O}_{q\times n}^{\times(l)}\\ E_{l}\\ \mathbb{O}_{q\times n}^{\times(r-l)}\\ \mathbb{O}_{m\times n}^{\times(l)}\\ \sqrt{\frac{1}{r+1}R}\,T(P_{\text{max}})\\ \mathbb{O}_{m\times n}^{\times(r-l)}\end{bmatrix}\in\mathbb{R}^{[(q+m)(r+1)]\times n},\,0\leq l\leq r.$
Since the system $\left[E_{0},E_{1},\dotsb,E_{r};A,C_{1},\dotsb,C_{r}\right]$ is stochastically detectable, there exists $\varTheta\in\mathbb{R}^{n\times q}$ such that $\left(A+\varTheta E_{0},C_{1}+\varTheta E_{1},\dotsb,C_{r}+\varTheta E_{r}\right)$ is mean-square stable. Let $\hat{\varTheta}\in\mathbb{R}^{n\times[(q+m)(r+1)]}$ and $\hat{\varTheta}$=
$\begin{bmatrix}\varTheta^{\times(r+1)}&-B\sqrt{(r+1)R^{-1}}&-D_{1}\sqrt{(r+1)R^{-1}}&\dotsb&-D_{r}\sqrt{(r+1)R^{-1}}\end{bmatrix}$,
then the system
$\left(A+BT(P_{\text{max}})+\hat{\varTheta}\hat{E}_{0},C_{1}+D_{1}T(P_{\text{max}})+\hat{\varTheta}\hat{E}_{1},\dotsb,C_{r}+D_{r}T(P_{\text{max}})+\hat{\varTheta}\hat{E}_{r}\right)$
is mean-square stable.
By Lemma 3.5, we get the system
$(A+BT(P_{\text{max}}),\ C_{1}+D_{1}T(P_{\text{max}}),\dotsb,C_{r}+D_{r}T(P_{\text{max}}))$
is mean-square stable. Hence, $P_{\text{max}}$ is a stabilizing solution of ARE (13). By Theorem 5.6.5 in [Dragan2013book], the stabilizing solution is unique, so ARE (13) has a unique stabilizing solution $P\in\overline{\mathbb{S}}_{+}^{n}$ such that $R+\sum_{l=1}^{r}D_{l}^{\top}PD_{l}\succ 0$. ∎
3.3 Convergence Analysis of Iterative Sequence
As a foundation for the convergence analysis, we define two linear operators associated with the interrelated iterative steps. Lemma 3.8 clarifies their essential properties, which are crucial because they collectively ensure the boundedness of the iteration sequence. This guarantee of boundedness is then directly utilized to prove convergence in Theorem 3.9.
For each triple $(P,Z,L)$ satisfying $P\in\mathrm{Dom}\,\mathcal{G}$, $P+Z\in\mathrm{Dom}\,\mathcal{G}$ and $L\in\mathcal{A}$, the linear operators $\mathcal{L}^{*}_{P}$ and $\mathcal{L}^{*}_{P+Z}$ are defined as follows:
$\mathcal{L}^{*}_{P}(Y)=Y(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}L)+(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}L)^{\top}Y+$ (14)
$\sum_{l=1}^{r}(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}L)^{\top}Y(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}L),\,\forall\,Y\in\mathbb{S}^{n};$
$\mathcal{L}^{*}_{P+Z}(Y)=Y(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L)+(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L)^{\top}Y+$ (15)
$\sum_{l=1}^{r}(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P+Z)+D_{l,2}L)^{\top}Y(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P+Z)+D_{l,2}L),\,\forall\,Y\in\mathbb{S}^{n},$
where $A_{(0)},C_{l,(0)}(1\leq l\leq r)$ are defined in (9) and
$\begin{bmatrix}\hat{K}_{1}(P)\\ \hat{K}_{2}(P)\end{bmatrix}=\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}-\begin{bmatrix}K_{1}(0)\\ K_{2}(0)\end{bmatrix},\quad\begin{bmatrix}K_{1}(P)\\ K_{2}(P)\end{bmatrix}=-R(P)^{-1}\begin{bmatrix}S_{1}(P)\\ S_{2}(P)\end{bmatrix}$. (16)
Lemma 3.8.
Assume $0\in\mathrm{Dom}\,\mathcal{G}$ and that there exists an $L\in\mathcal{A}$. Let $P,Z\in\mathbb{S}^{n}$ with the following properties:
  • $P,P+Z\in\mathrm{Dom}\,\mathcal{G}$.
  • $P$ and $Z$ satisfy:
    $0=\mathcal{G}(P)+Z(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}\hat{K}_{2}(P))+(A_{(0)}+B_{1}\hat{K}_{1}(P)+B_{2}\hat{K}_{2}(P))^{\top}Z+$ (17)
    $\sum_{l=1}^{r}(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))$
    $-[B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))]^{\top}R_{22}(P+Z)^{-1}$
    $\quad\times[B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))].$
If $\mathcal{L}^{*}_{P}$, $\mathcal{L}^{*}_{P+Z}$ are linear operators associated with the triple $(P,Z,L)$ via (14) and (15) and $\tilde{P}_{L}$ is the stabilizing solution of the ARE (5) associated with $L$, then the following assertions hold:
(i). If   $\operatorname{Spec}\,\mathcal{L}^{*}_{P}\subset\mathbb{C}^{-}$, then $\tilde{P}_{L}\succeq P+Z$.
(ii). If   $\tilde{P}_{L}\succeq P+Z$, then $\operatorname{Spec}\,\mathcal{L}^{*}_{P+Z}\subset\mathbb{C}^{-}$.
Proof.
First, we prove the (i). From Proposition 3.2 and Proposition 3.3, we can obtain:
$\mathcal{G}(P+Z)=$ $\mathcal{L}^{*}_{P}(P+Z)+Q_{L}+\hat{K}^{\top}_{1}(P)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P)+\hat{K}_{1}^{\top}(P)R_{11}\hat{K}_{1}(P)$
$-(J(P)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P)-\hat{K}(P+Z)),$
where $J(P)=\begin{bmatrix}\hat{K}_{1}(P)\\ L\end{bmatrix}$, $\mathcal{L}^{*}_{P}$ is defined in (14), $\hat{K}_{1}(P)$ and $\hat{K}(P)$ is defined in (16).
Combining (17) and Proposition 3.3, we find that $P+Z$ solves the following ARE:
$\mathcal{L}^{*}_{P}(P+Z)+Q_{L}+\hat{K}^{\top}_{1}(P)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P)+\hat{K}_{1}^{\top}(P)R_{11}\hat{K}_{1}(P)$ (18)
$-(J(P)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P)-\hat{K}(P+Z))$
$+(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}$
$\quad\times\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))=0,$
where $\hat{N}_{1}(P,Z)=B_{1}^{\top}Z+\sum_{l=1}^{r}D_{l,1}^{\top}Z(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))$, $\hat{N}_{2}(P,Z)=B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}Z\\ (C_{l,(0)}+D_{l,1}\hat{K}_{1}(P)+D_{l,2}\hat{K}_{2}(P))$ and $\mathbb{R}_{22}^{\sharp}(P+Z)$ is defined in (10). Moreover,
$(J(P)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P)-\hat{K}(P+Z))=$ (19)
$(\hat{K}_{1}(P)-\hat{K}_{1}(P+Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)(\hat{K}_{1}(P)-\hat{K}_{1}(P+Z))+H^{\top}_{1}H_{1}=$
$H^{\top}_{1}H_{1}+(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}$
$\quad\times(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z)),$
where $H_{1}=\begin{bmatrix}R_{22}(P+Z)^{-\frac{1}{2}}R_{21}(P+Z)&R_{22}(P+Z)^{\frac{1}{2}}\end{bmatrix}(J(P)-\hat{K}(P+Z))$.
Next, by virtue of Proposition 3.2, $\tilde{P}_{L}$ can be rewritten as:
$\mathcal{L}^{*}_{P}(\tilde{P}_{L})+Q_{L}+\hat{K}^{\top}_{1}(P)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P)+\hat{K}_{1}^{\top}(P)R_{11}\hat{K}_{1}(P)$ (20)
$-(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L}))=0,$
where $K_{L}(\tilde{P}_{L})=-(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}(B^{\top}_{1}\tilde{P}_{L}+\sum_{l=1}^{r}D^{\top}_{l,1}\tilde{P}_{L}C_{l}+R_{12}L)$.
Subtracting (18) from (20) and combining it with (19) yields:
$\mathcal{L}^{*}_{P}(\tilde{P}_{L}-P-Z)+H^{\top}H,$
$H^{\top}H=H_{1}^{\top}H_{1}-(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P)-K_{L}(\tilde{P}_{L})).$
Since the $\operatorname{Spec}\,\mathcal{L}^{*}_{P}\subset\mathbb{C}^{-}$ and $H^{\top}H\succeq 0$, by a similarly proof of (a) $\Rightarrow$ (b) in Lemma 3.5, we have $\tilde{P}_{L}\succeq P+Z$. This completes the proof of (i).
We now prove (ii). Similarly, applying Proposition 3.2, Proposition 3.3 and combining (17), we find that $P+Z$ solves the following ARE:
$\mathcal{L}^{*}_{P+Z}(P+Z)+\hat{K}^{\top}_{1}(P+Z)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P+Z)+\hat{K}_{1}^{\top}(P+Z)R_{11}\hat{K}_{1}(P+Z)$ (21)
$+Q_{L}-(J(P+Z)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P+Z)-\hat{K}(P+Z))$
$+(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}$
$\quad\times(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))=0.$
Additionally, by virtue of Proposition 3.2, $\tilde{P}_{L}$ can be rewritten as:
$\mathcal{L}^{*}_{P+Z}(\tilde{P}_{L})+\hat{K}^{\top}_{1}(P+Z)R_{12}L+L^{\top}R_{21}\hat{K}_{1}(P+Z)+\hat{K}_{1}^{\top}(P+Z)R_{11}\hat{K}_{1}(P+Z)+$ (22)
$Q_{L}-(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))=0,$
and $(J(P+Z)-\hat{K}(P+Z))^{\top}R(P+Z)(J(P+Z)-\hat{K}(P+Z))=$
$\begin{aligned} (\hat{K}_{1}(P+Z)-\hat{K}_{1}(P+Z))^{\top}\mathbb{R}_{22}^{\sharp}(P+Z)(\hat{K}_{1}(P+Z)-\hat{K}_{1}(P+Z))+H^{\top}_{2}H_{2}\end{aligned},$ (23)
where $H_{2}=\begin{bmatrix}R_{22}(P+Z)^{-\frac{1}{2}}R_{21}(P+Z)&R_{22}(P+Z)^{\frac{1}{2}}\end{bmatrix}(J(P+Z)-\hat{K}(P+Z))$.
Subtracting (21) from (22) and combining it with (23) yields:
$\mathcal{L}^{*}_{P+Z}(\tilde{P}_{L}-P-Z)+H_{2}^{\top}H_{2}-(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))^{\top}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})$
$\times(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))-(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))^{\top}$
$\times\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))=0.$
Let $\Delta=\tilde{P}_{L}-P-Z$, we have: $\mathcal{L}^{*}_{P+Z}\Delta+\sum_{l=0}^{r}\hat{E}_{l}^{\top}\hat{E}_{l}=0,\Delta\succeq 0$, where $\hat{E}_{l}\in\mathbb{R}^{[m_{2}+m_{1}(r+2)]\times n}(0\leq l\leq r)$,
$\hat{E}_{l}=\begin{bmatrix}\mathbb{O}_{m_{1}\times n}^{\times(l)}\\ \sqrt{-\frac{1}{r+1}(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})}(\hat{K}_{1}(P+Z)-K_{L}(\tilde{P}_{L}))\\ \mathbb{O}_{m_{1}\times n}^{\times(r-l)}\\ \frac{1}{\sqrt{r+1}}H_{2}\\ \sqrt{-\frac{1}{r+1}\mathbb{R}_{22}^{\sharp}(P+Z)^{-1}}(\hat{N}_{1}(P,Z)-R_{12}(P+Z)R_{22}(P+Z)^{-1}\hat{N}_{2}(P,Z))\end{bmatrix}$.
Choose $\hat{\varTheta}\in\mathbb{R}^{n\times[m_{2}+m_{1}(r+2)]}$, and
$\hat{\varTheta}=\begin{bmatrix}-\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}}^{\top}B_{1}^{\top}\\ -\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l,1})^{-1}}^{\top}D_{1,1}^{\top}\\ \dotsb\\ -\sqrt{-(r+1)(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}\tilde{P}_{L}D_{l})^{-1}}^{\top}D_{r,1}^{\top}\\ \mathbb{O}_{n\times m_{2}}^{\top}\\ \mathbb{O}_{n\times m_{1}}^{\top}\end{bmatrix}^{\top}$.
We obtain that $(A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L+\hat{\varTheta}\hat{E}_{0},C_{1,(0)}+D_{1,1}\hat{K}_{1}(P+Z)+D_{1,2}L+\hat{\varTheta}\hat{E}_{1},\dotsb,C_{r,(0)}+D_{r,1}\hat{K}_{1}(P+Z)+D_{r,2}L+\hat{\varTheta}\hat{E}_{r})$ associated with the system $(A_{L}+B_{1}K_{L}(\tilde{P}_{L}),C_{1L}+D_{1,1}K_{L}(\tilde{P}_{L}),\dotsb,C_{rL}+D_{r,1}K_{L}(\tilde{P}_{L}))$ is mean-square stable.
This implies the system $\hat{E}_{0},\hat{E}_{1},\dotsb,\hat{E}_{r};A_{(0)}+B_{1}\hat{K}_{1}(P+Z)+B_{2}L,C_{1,(0)}+D_{1,1}\hat{K}_{1}(P+Z)+D_{1,2}L,\dotsb,\\ C_{r,(0)}+D_{r,1}\hat{K}_{1}(P+Z)+D_{r,2}L]$ is stochastically detectable. By Lemma 3.5, we have $\operatorname{Spec}\,\mathcal{L}^{*}_{P+Z}\subset\mathbb{C}^{-}$. Thus, the proof of (ii). is completed. ∎
Before presenting the next theorem, we first introduce two families of linear operators: $\mathcal{L}^{*}_{P^{(k)}}$ and $\mathcal{L}^{(k,k+1)}$ ($k=0,1,2,\ldots$), which are associated with the iteration sequences $\{Z^{(k)}\}_{k\geq 0}$ and $\{P^{(k)}\}_{k\geq 0}$ defined in (2.3).
$\mathcal{L}^{*}_{P^{(k)}}(Y)=\,Y(A_{(0)}+B_{1}\hat{K}_{1}(P^{(k)})+B_{2}L)+(A_{(0)}+B_{1}\hat{K}_{1}(P^{(k)})+B_{2}L)^{\top}Y+$ (24)
$\sum_{l=1}^{r}(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P^{(k)})+D_{l,2}L)^{\top}Y(C_{l,(0)}+D_{l,1}\hat{K}_{1}(P^{(k)})+D_{l,2}L),\forall\,Y\in\mathbb{S}^{n},$
$\mathcal{L}^{(k,k+1)}(Y)=\,Y\left(A_{(k)}+B_{2}T_{(k+1)}\right)+\left(A_{(k)}+B_{2}T_{(k+1)}\right)^{\top}Y$ (25)
$+\sum_{l=1}^{r}\left(C_{l,(k)}+D_{l,2}T_{(k+1)}\right)^{\top}Y\left(C_{l,(k)}+D_{l,2}T_{(k+1)}\right),\forall\,Y\in\mathbb{S}^{n},$
where $T_{(k+1)}=-R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}$, $A_{(k)},C_{l,(k)}(1\leq l\leq r)$ and $N_{2,(k)}$ are defined in (9) and (10).
Theorem 3.9.
Assume the following conditions hold:
  • $R_{22}\succ 0$ and there exists an $L\in\mathcal{A}$.
  • There exist matrices $E_{l,(0)}(0\leq l\leq r)$ such that $\sum_{l=0}^{r}E_{l,(0)}^{\top}E_{l,(0)}=Q-S^{\top}(0)R(0)^{-1}S(0)$ and the system
    $\left[E_{0,(0)},E_{1,(0)},\dotsb,E_{r,(0)};A_{(0)},C_{1,(0)},\dotsb,C_{r,(0)}\right]$
    is stochastically detectable.
Then, we have:
(i). The sequences $\{Z^{(k)}\}_{k\geq 0}$, $\{P^{(k)}\}_{k\geq 0}$ are well defined by (6) (7) (8) , and for each $k=0,1,2\ldots$ the following items are fulfilled:
  1. $a_{k}$.
    $\operatorname{Spec}\,\mathcal{L}^{(k,k+1)}\subset\mathbb{C}^{-}$;
  2. $b_{k}$.
    Let $L\in\mathcal{A}$ be arbitrary but fixed and $\tilde{P}_{L}$ be the stabilizing solution to the corresponding ARE (5), then $\tilde{P}_{L}\succeq P^{(k)}+Z^{(k)}$ and $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(k+1)}}\subset\mathbb{C}^{-}$;
  3. $c_{k}$.
    $P^{(k)},P^{(k)}+Z^{(k)}\in\mathrm{Dom}\,\mathcal{G}$ and $\mathcal{G}(P^{(k)}+Z^{(k)})=$
    $-\left(N_{1,(k)}-R_{12}(P^{(k)}+Z^{(k)})R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}\right)^{\top}$
    $\times\mathbb{R}_{22}^{\sharp}(P^{(k)}+Z^{(k)})^{-1}$
    $\times\left(N_{1,(k)}-R_{12}(P^{(k)}+Z^{(k)})R_{22}(P^{(k)}+Z^{(k)})^{-1}N_{2,(k)}\right),$
    where $\mathbb{R}_{22}^{\sharp}(P^{(k)}+Z^{(k)}),N_{1,(k)}$ and $N_{2,(k)}$ are defined in (10);
(ii). $\lim_{k\rightarrow\infty}P^{(k)}=\tilde{P}$, where $\tilde{P}$ is the stabilizing solution of the stochastic GTARE (3).
Proof.
Step 1. When $k=0$ , $P^{(0)}=0$ and $Z^{(0)}$ is the stabilizing solution to the following ARE:
$ZA_{(0)}+A_{(0)}^{\top}Z+\sum_{l=1}^{r}C_{l,(0)}^{\top}ZC_{l,(0)}+Q-S^{\top}(0)R(0)^{-1}S(0)-(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(0)})^{\top}$ (26)
$\quad\times(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,2})^{-1}(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(0)})=0.$
To show that $Z^{(0)}$ is well-defined, we first prove the existence of a unique stabilizing solution to the above ARE.
Given that the set $\mathcal{A}$ is nonempty, let $L\in\mathcal{A}$ be arbitrary but fixed and set $L^{(0)}=L$. Then we can find the system $(A_{(0)}+B_{2}L^{(0)},C_{1,(0)}+D_{1,2}L^{(0)},\dotsb,C_{r,(0)}+D_{r,2}L^{(0)})$ associated with the system $(A_{L},C_{1L},\dotsb,C_{rL})$ is mean-square stable. This means the system
$A_{(0)},C_{1,(0)},\dotsb,C_{r,(0)};B_{2},D_{1,2},\dotsb,D_{r,2}]$
is mean-square stabilizable. Furthermore, the assumptions $R_{22}\succ 0$ and $\sum_{l=0}^{r}E^{\top}_{l,(0)}E_{l,(0)}=Q-S^{\top}(0)R(0)^{-1}S(0)\\ \succeq 0$ ensures that
$0\in\bm{\Gamma}^{[A_{(0)},C_{1,(0)},\dotsb,C_{r,(0)};B_{2},D_{1,2},\dotsb,D_{r,2};\sum_{l=0}^{r}E^{\top}_{l,(0)}E_{l,(0)},R_{22}]}.$
And the system $E_{0,(0)},E_{1,(0)},\,\,\dotsb\,\,,E_{r,(0)};A_{(0)},C_{1,(0)},\,\,\dotsb\,\,,C_{r,(0)}]$ is stochastically detectable. By virtue of Proposition 3.7, the ARE (26) admits a unique stabilizing solution $Z^{(0)}$ such that $R_{22}(Z^{(0)})\succ 0$. Thus, $Z^{(0)}$ is well-defined as the unique stabilizing solution to the ARE (26). This further implies that $\operatorname{Spec}\,\mathcal{L}^{(0,1)}\subset\mathbb{C}^{-}$.
Let $\tilde{P}_{L}$ denote the stabilizing solution to the ARE (5) associated with $L$. Since the operator $\mathcal{L}^{*}_{P^{(0)}}$ is associated with the system $(A_{L},C_{1L},\dotsb,C_{rL})$ is mean-square stable, we have the $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(0)}}\subset\mathbb{C}^{-}$. From (i) and (ii) in Lemma 3.8, we get $\tilde{P}_{L}\succeq P^{(0)}+Z^{(0)}$ and $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(1)}}\subset\mathbb{C}^{-}$.
Combining $R_{22}(\tilde{P}_{L})\succeq R_{22}\succ 0$ and $R_{11}\preceq R_{11}(\tilde{P}_{L})\prec 0$, we have $P^{(0)}$ and $\tilde{P}_{L}\in\mathrm{Dom}\,\mathcal{G}$. Since $R_{22}(P^{(0)}+Z^{(0)})\succeq R_{22}\succ 0$ and $R_{11}(P^{(0)}+Z^{(0)})\prec R_{11}(\tilde{P}_{L})\prec 0$, we obtain that $P^{(0)}+Z^{(0)}\in\mathrm{Dom}\,\mathcal{G}$.
Substituting the ARE (26) into $\mathcal{G}(P^{(0)}+Z^{(0)})$ by using Proposition 3.3, then gives $\mathcal{G}(P^{(0)}+Z^{(0)})=(N_{1,(0)}-R_{12}(Z^{(0)})R_{22}(Z^{(0)})^{-1}N_{2,(0)})^{\top}R^{\sharp}_{22}(Z^{(0)})^{-1}(N_{1,(0)}-R_{12}(Z^{(0)})R_{22}(Z^{(0)})^{-1}N_{2,(0)})$.
This completes the proof of statements $a_{0}-c_{0}$.
Step 2. Assume k=h1k=h-1, $Z^{(h-1)}$ is well-defined and that $a_{h-1}-c_{h-1}$ hold. We shall prove that $Z^{(h)}$ is well-defined and that $a_{h}-c_{h}$ hold.
For $k=h$ , $Z^{(h)}$ satisfies the following ARE:
$ZA_{(h)}+A_{(h)}^{\top}Z+\sum_{l=1}^{r}C_{l,(h)}^{\top}ZC_{l,(h)}+\sum_{l=0}^{r}E^{\top}_{l,(h)}E_{l,(h)}-(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(h)})^{\top}$ (27)
$\quad\times(R_{22}(P^{(h)})+\sum_{l=1}^{r}D_{l,2}^{\top}ZD_{l,2})^{-1}(B_{2}^{\top}Z+\sum_{l=1}^{r}D_{l,2}^{\top}ZC_{l,(h)})=0,$
where $E_{l,(h)}\in\mathbb{R}^{[m_{1}(r+1)]\times n}(0\leq l\leq r)$, and
$E_{l,(h)}=\begin{bmatrix}\mathbb{O}_{m_{1}\times n}^{\times(l)}\\ \sqrt{-\frac{1}{r+1}\mathbb{R}_{22}^{\sharp}\left(P^{(h)}\right)^{-1}}\left(N_{1,(h-1)}-R_{12}(P^{(h)})R_{22}(P^{(h)})^{-1}N_{2,(h-1)}\right)\\ \mathbb{O}_{m_{1}\times n}^{\times(r-l)}\end{bmatrix}$.
We demonstrate that $Z^{(h)}$ is well-defined. By the Proposition 3.7, proving that $Z^{(h)}$ is well-defined is equivalent to proving that the system
$\left[A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)};B_{2},D_{1,2},\dotsb,D_{r,2}\right]$
is mean-square stabilizable and the system
$\left[E_{0,(h)},E_{1,(h)},\dotsb,E_{r,(h)};A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)}\right]$
is stochastically detectable.
To prove the mean-square stabilizability of the system
$A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)};B_{2},D_{1,2},\dotsb,D_{r,2}],$
it suffices to set $L^{(h)}=L-\hat{K}_{2}(P^{(h)})$, where $\hat{K}_{2}(P)$ is defined in (16). We can then identify the system
$(A_{(h)}+B_{2}L^{(h)},C_{1,(h)}+D_{1,2}L^{(h)},\dotsb,C_{r,(h)}+D_{r,2}L^{(h)})$
associated with the operator $\mathcal{L}^{*}_{P^{(h)}}$. Given that the condition $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(h)}}\subset\mathbb{C}^{-}$, we thereby prove that the system
$\left[A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)};B_{2},D_{1,2},\dotsb,D_{r,2}\right]$
is mean-square stabilizable. Let
$\tilde{\varTheta}^{(h)}=\begin{bmatrix}\tilde{\varTheta}_{0}^{(h)}&\dotsb&\tilde{\varTheta}_{r}^{(h)}\end{bmatrix}\in\mathbb{R}^{n\times[m_{1}(r+1)]},$
where
$\tilde{\varTheta}_{0}^{(h)}=-(B_{1}-B_{2}R_{22}(P^{(h)})^{-1}R_{21}(P^{(h)}))\sqrt{-(r+1)\mathbb{R}_{22}^{\sharp}\left(P^{(h)}\right)^{-1}}$
and
$\tilde{\varTheta}_{l}^{(h)}=-(D_{l1}-D_{l2}R_{22}(P^{(h)})^{-1}R_{21}(P^{(h)}))\sqrt{-(r+1)\mathbb{R}_{22}^{\sharp}\left(P^{(h)}\right)^{-1}},1\leq l\leq r.$
Then we can find the system
$(A_{(h)}+\tilde{\varTheta}^{(h)}E_{0,(h)},C_{1,(h)}+\tilde{\varTheta}^{(h)}E_{1,(h)},\dotsb,C_{r,(h)}+\tilde{\varTheta}^{(h)}E_{r,(h)})$
associated with $\mathcal{L}^{(h-1,h)}$ is mean-square stable. This means the system
$E_{0,(h)},E_{1,(h)},\dotsb,E_{r,(h)};A_{(h)},C_{1,(h)},\dotsb,C_{r,(h)}]$
is stochastically detectable. Similarly, from Proposition 3.7, the ARE (27) admits a unique stabilizing solution $Z^{(h)}$ such that $R_{22}(P^{(h)}+Z^{(h)})\succ 0$ and we also get $\operatorname{Spec}\,\mathcal{L}^{(h,h+1)}\subset\mathbb{C}^{-}$.
Since $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(h)}}\subset\mathbb{C}^{-}$, from (i) and (ii) in Lemma 3.8, we get $\tilde{P}_{L}\succeq P^{(h)}+Z^{(h)}$ and $\operatorname{Spec}\,\mathcal{L}^{*}_{P^{(h+1)}}\subset\mathbb{C}^{-}$. Similarly, combining $R_{22}(P^{(h)}+Z^{(h)})\succ 0$ and $R_{11}(P^{(h)}+Z^{(h)})\preceq R_{11}(\tilde{P}_{L})\prec 0$, we obtain that $P^{(h)}+Z^{(h)}\in\mathrm{Dom}\,\mathcal{G}$. Substituting (27) into $\mathcal{G}(P^{(h)}+Z^{(h)})$ by using Proposition 3.3, then gives $\mathcal{G}(P^{(h)}+Z^{(h)})=(N_{1,(h)}-R_{12}(P^{(h)}+Z^{(h)})R_{22}(P^{(h)}+Z^{(h)})^{-1}N_{2,(h)})^{\top}R^{\sharp}_{22}(P^{(h)}+Z^{(h)})^{-1}(N_{1,(h)}-R_{12}(P^{(h)}+Z^{(h)})R_{22}(P^{(h)}+Z^{(h)})^{-1}N_{2,(h)})$.
Thus, we have proved the statements $a_{h}-c_{h}$. By induction, we conclude that for any $k$, $Z^{(k)}$ is well-defined and $a_{k}-c_{k}$ hold.
Step 3. In this recursive process, the sequence $\{P^{(k)}\}_{k\geq 0}$ is monotonically non-decreasing and bounded above, so the sequence $\{P^{(k)}\}_{k\geq 0}$ is convergent and
$\lim_{h\rightarrow\infty}\sum_{s=h+1}^{\infty}Z^{(s)}=0,\lim_{k\rightarrow\infty}Z^{(k)}=0.$
Set $P^{*}=\lim_{k\rightarrow\infty}P^{(k)}$, given that
$\mathcal{G}(P^{*})=\mathcal{G}(\lim_{k\rightarrow\infty}P^{(k)})=\mathcal{G}\left(\sum_{s=0}^{\infty}Z^{(s)}\right)=$
$\sum_{s=h+1}^{\infty}Z^{(s)}A_{(h)}+\sum_{s=h+1}^{\infty}A_{(h)}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}C_{l,(h)}^{\top}Z^{(s)}C_{l,(h)}$
$-\left(\sum_{s=h+1}^{\infty}B_{2}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(s)}C_{l,(h)}\right)^{\top}\left(R_{22}+\sum_{l=1}^{r}D_{l,2}^{\top}P^{*}D_{l,2}\right)^{-1}$
$\quad\times\left(\sum_{s=h+1}^{\infty}B_{2}^{\top}Z^{(s)}+\sum_{s=h+1}^{\infty}\sum_{l=1}^{r}D_{l,2}^{\top}Z^{(s)}C_{l,(h)}\right)+\mathcal{G}\left(\sum_{s=0}^{h}Z^{(s)}\right),\forall h=0,1,2,\dotsb.$
Since $\lim_{h\rightarrow\infty}\sum_{s=h+1}^{\infty}Z^{(s)}=0$ and $\lim_{h\rightarrow\infty}\mathcal{G}\left(\sum_{s=0}^{h}Z^{(s)}\right)=$
$\lim_{h\rightarrow\infty}(N_{1,(h)}-R_{12}(P^{(h+1)})R_{22}(P^{(h+1)})^{-1}N_{2,(h)})^{\top}R^{\sharp}_{22}(P^{(h+1)})^{-1}$
$\times(N_{1,(h)}-R_{12}(P^{(h+1)})R_{22}(P^{(h+1)})^{-1}N_{2,(h)})=0,$
it follows that $\mathcal{G}(P^{*})=0$ and $P^{*}\succeq 0$ is the solution to the stochastic GTARE (3).
[…]
By Proposition 3.3, expanding $\mathcal{G}(P^{*})$ via $P^{*}=(P^{*}-\tilde{P})+\tilde{P}$ yields:
[…]
$)\displaystyle+\sum_{l=1}^{r}(C_{l}+D_{l,1}K_{1}(\tilde{P})+D_{l,2}K_{2}(\tilde{P}))^{\top}(P^{*}-\tilde{P})(C_{l}+D_{l,1}K_{1}(\tilde{P})+D_{l,2}K_{2}(\tilde{P}$))
[…]
From Proposition 3.7, we have $P^{*}\succeq\tilde{P}$. Combining this with $\tilde{P}_{\tilde{L}}=\tilde{P}$, we obtain $\tilde{P}\preceq P^{*}\preceq\tilde{P}_{\tilde{L}}=\tilde{P}$. Therefore, $P^{*}=\tilde{P}$ is the unique stabilizing solution to the stochastic GTARE (3). ∎
4 Numerical Example
To verify the effectiveness of the proposed computational method, we present a specific numerical example and randomly generate the system parameters as follows:
$A$ $=\begin{bmatrix}0.951718&-0.270753&-0.399002\\ -0.402652&-0.332890&-0.211269\\ 0.865505&0.242714&0.051286\end{bmatrix}$ $Q$ $=\begin{bmatrix}3.030930&-1.047214&-0.749947\\ -1.047214&2.194723&1.675449\\ -0.749947&1.675449&3.333918\end{bmatrix}$
$C_{1}$ $=\begin{bmatrix}0.428394&-0.231063&1.819921\\ 1.636217&-0.730599&1.212696\\ -0.187034&-0.538025&1.193243\end{bmatrix}$ $C_{2}$ $=\begin{bmatrix}0.160803&-1.563720&-1.122872\\ -0.855569&-0.072125&0.710778\\ 0.091879&0.884612&0.977433\end{bmatrix}$
$B_{1}$ $=\begin{bmatrix}0.325576&-0.086447&-0.227595\\ -0.281475&0.347989&-0.397287\\ -0.079578&-0.101469&-0.095411\end{bmatrix}$ $B_{2}$ $=\begin{bmatrix}1.102509&0.029449&0.441306\\ 0.188708&1.055548&0.363029\\ 0.257869&0.036254&1.116683\end{bmatrix}$
$D_{11}$ $=\begin{bmatrix}0.007102&0.008817&0.004927\\ 0.006974&0.000950&0.001090\\ 0.009301&0.004565&0.001537\end{bmatrix}$ $D_{12}$ $=\begin{bmatrix}0.005681&0.002687&0.000159\\ 0.000953&0.009114&0.008537\\ 0.007980&0.009017&0.006756\end{bmatrix}$
$D_{21}$ $=\begin{bmatrix}0.009843&0.001642&0.003074\\ 0.002716&0.001324&0.004221\\ 0.008972&0.003174&0.003310\end{bmatrix}$ $D_{22}$ $=\begin{bmatrix}0.000377&0.005941&0.009387\\ 0.003089&0.006808&0.007371\\ 0.005642&0.006324&0.007449\end{bmatrix}$
$R_{11}$ $=\begin{bmatrix}-1.658649&-0.888904&-0.707647\\ -0.888904&-1.960809&-0.586128\\ -0.707647&-0.586128&-1.471209\end{bmatrix}$ $R_{12}$ $=\begin{bmatrix}0.319075&0.370440&0.509807\\ 0.212195&0.132083&0.857426\\ 0.313025&0.677299&0.000364\end{bmatrix}$
$R_{21}$ $=\begin{bmatrix}0.319075&0.212195&0.313025\\ 0.370440&0.132083&0.677299\\ 0.509807&0.857426&0.000364\end{bmatrix}$ $R_{22}$ $=\begin{bmatrix}1.078023&0.529765&0.411900\\ 0.529765&0.908261&0.312832\\ 0.411900&0.312832&1.135871\end{bmatrix}$
$S_{1}$ $=\begin{bmatrix}0.779145&0.430019&0.484125\\ 0.792295&0.346524&0.258780\\ 0.064140&0.841675&0.116926\end{bmatrix}$ $S_{2}$ $=\begin{bmatrix}0.616144&0.362532&0.803394\\ 0.351730&0.980676&0.676821\\ 0.513458&0.405686&0.268156\end{bmatrix}$
After 13 external iterations, the iterative sequence converges to the stabilizing solution of the problem within the specified error tolerance. The norm of the equation residual is $4.9409\times 10^{-6}$. The specific values of the stabilizing solution to the stochastic GTARE (3) (retained to five decimal places) are presented as follows:
$\begin{bmatrix}4.49040&0.33056&-1.00210\\ 0.33056&4.89533&-0.78154\\ -1.00210&-0.78154&8.61632\end{bmatrix}$.
Figure 1: The figure shows the real parts of three eigenvalues of $Z^{(k-1)}\,(k=1,\dotsb,13)$ evolving with external iterations.
Figure 2: The figure shows the real parts of five eigenvalues of $M_{(k-1)}(k=1,\dotsb,13)$ evolving with external iterations.
Figure 3: This simulation diagram clearly illustrates the dynamic process of the system state under the balanced control strategies of two players.
Fig. 1 and Fig. 2 illustrate the evolution of the real parts of eigenvalues across different matrices over external iterations $(k=1,\dotsb,13)$. Both figures reveal a consistent trend: the real parts of the eigenvalues of $Z^{(k)}$ and $M_{(k)}$ respectively decrease with iterations, approaching 0 by the sixth iteration and continuing to decline thereafter. Notably, the eigenvalues of $M_{(1)}$ undergo a drastic change, attributed to one-time disturbances from the opposing player. These collectively reflect an interesting characteristic: when the set $\mathcal{A}$ is non-empty, Player 2 initiates from a dominant position and applies controls to stabilize the system. Meanwhile, Player 1 is dissatisfied with the current values of the performance functional. In response, Player 2 makes tentative concessions by continuously granting Player 1 certain advantages, leading to a gradual reduction in Player 1’s dissatisfaction. Through such continuous adjustments, the performance functional gradually approaches its equilibrium values.
Fig. 3 illustrates the dynamic evolution of the system’s three state variables ($x_{1}$, $x_{2}$, $x_{3}$) starting from a randomly generated initial state. Initially, the system displays pronounced unstable dynamics, evidenced by the rapid and intense transients of all states—with $x_{3}$ exhibiting a particularly large initial deviation. This unstable trend is quickly suppressed, however, and the states converge asymptotically to the origin. The control input diagrams in the middle and lower panels explain this phenomenon: the controls exerted by the two players involve strong, oppositely directed antagonistic injections. Within the stability injection set $\mathcal{K}$, the two players pursue mutually acceptable equilibrium control strategies to ensure the system’s equilibrium across all directions. While their control injection are antagonistic, their interaction yields a combined effect that counteracts the system’s inherent unstable dynamics, collectively driving the states toward equilibrium. Notably, despite the zero-sum game framework, the two parties demonstrate a cooperative tendency in maintaining system stability. These observations align with the theoretical findings presented in Remark 2.6.3 of [Sun2020_book].
Together, these figures verify the effectiveness of our computational method and indirectly reveal the mechanism by which the method iteratively finds the equilibrium solution.